### v3

# A von Neumann algebraic approach to Quantum Theory on curved spacetime

Abstract.
By extending the method developed in our recent paper [31] we present the AQFT framework in terms of von Neumann algebras. In particular, this approach allows for a locally covariant categorical description of AQFT which moreover satisfies the additivity property and provides a natural and intrinsic framework for a description of entanglement. Turning to dynamical aspects of QFT we show that Killing local flows may be lifted to the algebraic setting in curved space-time. Furthermore, conditions under which quantum Lie derivatives of such local flows exist are provided. The central question that then emerges is how such quantum local flows might be described in interesting representations. We show that quasi-free representations of Weyl algebras fit the presented framework perfectly. Finally, the problem of enlarging the set of observables is discussed. We point out the usefulness of Orlicz space techniques to encompass unbounded field operators. In particular, a well-defined framework within which one can manipulate such operators is necessary for the correct presentation of (semiclassical) Einstein’s equation.
Key words and phrases:
local quantum field theory, non-commutative measure and integration, derivations, non-commutative geometry
2010 Mathematics Subject Classification:
81T05, 46L51, 47L90 (primary); 46E30, 58A05,46L52, (secondary)
This manuscript has no associated data.
The authors whose names are listed above declare that they have no conflicts of interest regarding the subject matter or materials discussed in this manuscript.
1. Introduction
It is generally accepted that observables and states are basic concepts in the description of a physical system. The description of a quantum system which emphasises operators representing observables leads to the algebraic approach to Quantum Theory. For a motivation of such approach we refer the reader to the book of Emch [17]. Then let us note that in Quantum Field Theory (QFT for short) there are two axiomatic programmes for QFT. The one which focuses on local observables and puts emphasis on operators is called Algebraic Quantum Field Theory (AQFT for short). The second one, more traditional, focuses on fields and Wightman functions and leads to the so called Gårding-Wightman programme, for details see the Haag book [26].
The basic feature of AQFT is its emphasis on the description of algebraic relations between the primary objects of the theory. In other words, fields and observables are treated as purely algebraic (and abstract) objects. On the other hand, when examining a specific system in a specific situation, in addition to algebraic relations between objects, we must use the available knowledge about the system. This knowledge is described by the state of the system. Therefore, in the analysis of field systems, one should move to a representation that takes into account algebraic relations between the examined objects as well as the current state of the object (field). The implementation of such a program involves the use of the GNS (Gelfand-Naimark-Siegel) structure. As a result, we obtain a representation of the initial algebraic structure on a specific Hilbert space, with the considered representation reflecting the specific state of the described system.
Recalling that the result of a physical measurement of a specific observable is described by the value of the state of the system on that observable, reasonable topologies from the point of view of a physicist are weak topologies. Therefore, it is natural for an algebraic structure to be additionally equipped with a weak topology. Since von Neumann algebras satisfy the above conditions, they will be the focus of the following analysis. In concluding our general remarks about representations, let us add that, on moving to (faithful) representations, the set of states in the representation (folium) is a dense set in the set of states of the initial algebra, see Section III.2 in [26]. Therefore, the transition to a representation is not in essence a limitation of the description.
In the introduction to [10] the authors rightly note that since in AQFT all structures are formulated in terms of local quantities, it is inappropriate to demand the a priori existence of a distinguished global Hilbert space of states. The authors also note that whichever category of operator algebras is used to model the theory needs to be a tensorial category. By basing the theory on abstract $C^{*}$-algebras rather than von Neumann algebras, the authors at the outset exclude any difficulties regarding Hilbert spaces since none are required. As pointed out on pages 8 and 9 of the related paper [22], this choice also circumvents problems related to generalizing the topological field theory construction to the case of infinite dimensional Hilbert spaces. Selecting $C^{*}$-algebras as the basic object may help to avoid some of the complexities involved incorporating Hilbert spaces into the picture but on the downside $C^{*}$-algebras allow for a menagerie of “canonical” tensor products. More seriously the exclusion of Hilbert spaces also excludes any chance of a refined theory of entanglement.
The important point to note here is the fact that $C^{*}$-algebras can be considered as a noncommutative analogue of the family of continuous functions while von Neumann algebras can be considered as a noncommutative counterpart of measurable functions. This fact places von Neumann algebras at the heart of noncommutative integration theory. In our works, [35, 33, 31], we proposed a description of large quantum systems (statistical physics, QFT) based on noncommutative analysis, i.e. on noncommutative integration theory and noncommutative differentiation. Since as noted above (abstract) von Neumann algebras are at the heart of noncommutative integration theory, we shall therefore use abstract von Neumann algebras when engaged in algebraic quantization. We emphasize that the use of abstract von Neumann algebras is not bound in any way to the choice of any particular state. See section 3 for details. In fact as we shall see in section 7 good states which well illustrate emerge from the very geometry of the underlying Lorentzian manifold.
One very obvious advantage of the von Neumann algebraic framework is the uniqueness of the von Neumann algebra tensor product. So the problem of having to decide which tensor product to use never arises. But this raises the question of how the requirement of not demanding a distinguished a priori Hilbert space is to be satisfied. We will demonstrate that the theory of the standard form of von Neumann algebras defines a tensorial category which enables one to incorporate Hilbert spaces into the picture at a local level. As such it fulfils the criterion of not demanding the a priori existence of a distinguished global Hilbert space of states, but nevertheless leaves room for a theory of entanglement at a local level.
In concluding this section we draw attention to some recent papers on this general topic written from a von Neumann algebra perspective. Specifically the following:
  1. (1)
    An algebra of observables for de Sitter space by Chandrasekaran, Longo, Penington, Witten [12]. Here the authors describe an algebra of observables for a static patch in de Sitter space, with operators gravitationally dressed to the worldline of an observer. The algebra is a von Neumann algebra of Type $II_{1}$. There is a natural notion of entropy for a state of such an algebra. In their paper the authors also point out the difficulty of isolating a concept of entropy for type III factors. As a contribution to this debate the authors of the present paper wish to point out that in one of their earlier papers [35, Definition 5.3], they identified a class of states who do admit a concept of entropy which proves to be a natural generalisation of von Neumann entropy.
  2. (2)
    Gravity and the crossed product by Witten[52]: In the context of quantum gravity Witten here uses crossed products to give meaning to the notion of the $1/N$ expansion of the ambient algebra. Passing to the crossed product then enables him to introduce a notion of entropy for this system. One of the main benefits of passing to the crossed product (as also explained by Witten) is that it will transform the modular automorphism group - which is not inner for type III algebras - to an inner automorphism. The generator of this automorphism, which is some sort of Hamiltonian, is then affiliated to this enlarged algebra. The value of Witten’s paper is that it shows that in many contexts crossed products find application in a natural way.
    • We pause to remind the reader that crossed products werre similarly used in [31] to introduce the notion of $L^{\cosh-1}$-regularity which is a regularity restriction ensuring that for a large number of examples the field operators of Minkowski space may in a natural way be embedded into the Orlicz space $L^{\cosh-1}({\mathcal{M}})$, which in turn may be viewed as a space which is home to “moment generating” observables.
  3. (3)
    Algebras and states in JT gravity by Penington, Witten [42]. The authors analyze the algebra of boundary observables in canonically quantised JT gravity with or without matter. In the absence of matter, this algebra is commutative, generated by the ADM Hamiltonian. After coupling to a bulk quantum field theory, it becomes a highly noncommutative algebra of type $II_{\infty}$ with a trivial center.
The above papers indicate that the von Neumann algebra approach still has strong support and that for at least for some models, it is to be preferred to the $C^{*}$-algebra approach. The above remarks should also indicate why in this article we choose to deal with AQFT based on von Neumann algebras.
The theme running throughout the paper is the suitability of von Neumann algebras for AQFT. One objection that has occasionally been raised in this regard is the fact that von Neumann algebras require an a priori Hilbert space which would seem to impose a random restriction on the mathematical formalism modelling AQFT. We address this concern is §3. We specifically show that standard forms of von Neumann algebras do in fact form a tensorial category, and that in this formalism each standard form comes equipped with its own unique “canonical” Hilbert space. In this formalism the associated Hilbert space is not some random unrelated structure imposed at the start, but a property of the system itself. Another significant feature of this category is that unlike the $C^{*}$ picture, it allows for the incorporation of entanglement into AQFT.
In part 1 of the paper (comprising sections 4-6) we investigate the role of (quantum) Killing vector fields in AQFT. We show that in important examples local flows of Killing vector fields may locally be lifted to the algebra level. We also provide very mild criteria under which infinitesimal generators of this lifted action exist as (locally) densely defined *-derivations – operators we may regard as quantum Lie derivatives. On Minkowski space Killing vector fields correspond to the action of the Poincaré group. Thus the philosophical significance of the achieved theory is that this local action of Killing vector fields at the algebraic level may be seen as a curved version of the action of the Poincare group.
In part 2 (sections 7-9) we consider local von Neumann algebras constructed from appropriate representations of Bosonic $\mathrm{CCR}$ algebras. Two questions are addressed here: Firstly an identification of those states which yield “good” representations of such algebras, and secondly a consideration of the extent to which the applications of Orlicz space geometry to AQFT, as demonstrated in [31] for the Minkowski space setting, can be carried over to the curved setting. In part 3 we finally discuss conclusions and options for further development.
2. Preliminaries
Throughout the pair $(M,g)$ will denote a Lorentzian manifold $M$ equipped with a Lorentzian metric $g$. Such manifolds will also sometimes be referred to as spacetimes. A particularly class of Lorentzian manifolds are the globally hyperbolic manifolds. There are several equivalent descriptions of global hyperbolicity. Two particularly elegant conditions for global hyperbolicity are that the spacetime admits a Cauchy hypersurface, or alternatively that $M$ is isometric to $\mathbb{R}\times S$ with metric $\beta dt^{2}+g_{t}$ where $\beta$ is a smooth positive function, $g_{t}$ is a Riemannian metric on $S$ depending smoothly on $t\in\mathbb{R}$ and each $\{t\}\times S$. We shall in this paper restrict attention to globally hyperbolic spacetimes. Given a globally hyperbolic spacetime $M$, the class of open relatively compact globally hyperbolic submanifolds $\mathcal{O}\subset M$ will be denoted by $\mathcal{K}(M,g)$. The precise content of what we mean by this statement is given in the discussion following definition 4.1.
As was the case in [31], we shall here be concerned with local algebras generated by field operators which are solutions of the Klein-Gordon equation. $C^{\infty}_{0}(M)$ will denote the space of smooth, real valued functions on $M$ which have compact support. Following Dimock [15], Brunetti, Fredehagen and Verch [9] and Bär, Ginoux and Pfäffle [3], we will on the manifold $M$ describe the $\mathrm{CCR}$ algebra of bosonic fields given by solutions of the Klein-Gordon equation. The global hyperbolicity of $M$ guarantees the existence of global fundamental solutions $\psi$ for the Klein-Gordon equation $(\Box+m^{2}+\xi R)\phi=0$, where $m\geq 0$ , $\xi\geq 0$ are constants, and $R$ is the scalar curvature of the metric on $M$. Given any vector bundle $E\to M$, we write $\mathscr{D}(M,E)$ for the space of compactly supported smooth sections in $E$. The solutions of the Klein-Gordon equation are of the form $G(f)$ ($f\in\mathscr{D}(M,E))$, where the operator $G$ is $G=G^{+}-G^{-}$ where $G^{+/-}$ are respectively the advanced/retarded Green’s operators $G^{+/-}:\mathscr{D}(M,E)\to C^{\infty}(M,E)$.
When $C^{*}$-algebras are in view we shall denote them by $\mathcal{A},\mathcal{B}$, with the symbols $\mathcal{M},\mathcal{N}$ being used to denote von Neumann algebras. Given a faithful normal semifinite weight $\nu$ on a von Neumann algebra, one may of course pass to the GNS-representation $(\pi_{\nu}:\mathcal{M}\to B(H_{\nu}))$ of $\mathcal{M}$} with respect to $\nu$. We will hereafter abbreviate faithful normal semifinite to just fns. With $\eta$ denoting the canonical embedding of the left-ideal $\mathfrak{n}_{\nu}=\{a\in\mathcal{M}\colon\nu(a^{*}a)<\infty\}$ into $H_{\nu}$, one may of course densely define an antilinear closable operator $S_{0}$ on $H_{\nu}$ by the prescription $\eta(a)\to\eta(a^{*})$, $a\in\mathcal{M}$. With $S$ denoting the closure of $S_{0}$, the operator $\Delta=S^{*}S$ is then referred to as the modular operator and the anti-unitary $J$ in the polar decomposition $S=J\Delta^{1/2}$ as modular conjugation. It is well known that the prescription $\sigma_{t}^{\nu}:a\to\pi_{\nu}^{-1}(\Delta^{it}\pi_{\nu}(a)\Delta^{-it})$ $a\in\mathcal{M}$ defines an automorphism group on $\mathcal{M}$} (the so called modular automorphism group) and $j_{\nu}:f\to JfJ$ ($J\in\pi_{\nu}(\mathcal{M})$) an anti-linear *-isomorphism from $\pi_{\nu}(\mathcal{M})$ to $\pi_{\nu}(\mathcal{M})^{\prime}$. On for the sake of ease of notation identifying $\mathcal{M}$ with $\pi_{\nu}(\mathcal{M})$, the natural positive cone $\mathscr{P}_{\nu}$ of $H_{\nu}$ is then defined to be the closure of the set $\{\eta{aj(a)}\colon a\in\mathcal{M}\}$. The content of Haagerup’s theorem regarding the standard form of a von Neumann algebra, is that the quintuple $(\mathcal{M},\pi_{\nu},H_{\nu},j_{\nu},\mathscr{P}_{\nu})$ uniquely identifies the GNS representation of $\mathcal{M}$ up to a spatial *-isomorphism.
Modular theory is also the gateway to the theory of quantum $L^{p}$ and Orlicz spaces, in that for general von Neumann algebras $\mathcal{M}$ one first needs to enlarge the algebra to the crossed product ${\mathfrak{M}}=\mathcal{M}\rtimes_{\nu}\mathbb{R}$. This algebra admits a canonical faithful normal semifinite trace which then allows us to further enlarge the algebra to the algebra of $\tau$-measurable operators $\widetilde{{\mathfrak{M}}}$ affiliated with $\mathfrak{M}$}. The actual construction of quantum $L^{p}$ and Orlicz spaces then happens inside this algebra; for brief details regarding this process we refer the reader to [31], and to [25] for a detailed account. We shall here have occasion to occasionally refer to the Orlicz space $L^{\cosh-1}({\mathcal{M}})$.
The focus of this paper is of course local algebras of Lorentzian spacetimes. These are collections of $C^{*}$ (alternatively von Neumann) algebras ${\mathcal{A}}(\mathcal{O})$ each corresponding to an element $\mathcal{O}$ of some regular class of submanifolds in such a way the entire collection indexed by the submanifolds forms what is known as a (weak) quasi-local algebra. The basic idea here is that each local algebra ${\mathcal{A}}(\mathcal{O})$ is a home for the observables of that region of spacetime. For the sake of the reader we record the definition of a quasi-local $C^{*}$-algebra as given in [3]. The indexing set $I$ needs to be a directed set satisfying a (weak) orthogonality relation. Specifically a set $I$ is a directed set satisfying an orthogonality relation if it is a partially ordered set also carrying a relation $\perp$ in such a way that
  1. (1)
    for all $\alpha,\beta\in I$ there exists some $\gamma\in I$ with $\gamma\geq\alpha$ and $\gamma\geq\beta$;
  2. (2)
    for every $\alpha\in I$ we can find some $\beta\in I$ with $\beta\perp\alpha$;
  3. (3)
    if $\alpha\leq\beta$ and $\beta\perp\gamma$ then also $\alpha\perp\gamma$;
  4. (4)
    if $\alpha\perp\beta$ and $\alpha\perp\gamma$, there exists $\delta\in I$ such that $\delta\geq\beta$, $\delta\geq\gamma$ and $\delta\perp\alpha$
If only conditions (1)-(3) are satisfied we refer to the orthogonality relation as a weak orthogonality relation. A quasi-local algebra (alt. weak quasi-local algebra) is a pair $({\mathcal{A}},\{{\mathcal{A}}_{\alpha}\}_{\alpha\in I})$ where $I$ is a directed set satisfying an orthogonality relation (alt. weak orthogonality relation) such that the following holds:
  • (i)
    ${\mathcal{A}}_{\alpha}\subset{\mathcal{A}}_{\beta}$ whenever $\alpha\leq\beta$;
  • (ii)
    the algebras ${\mathcal{A}}_{\alpha}$ all carry a common unit;
  • (iii)
    $\cup_{\alpha}{\mathcal{A}}_{\alpha}$ is norm-dense in $\mathcal{A}$};
  • (iv)
    if $\alpha\perp\beta$ then ${\mathcal{A}}_{\alpha},{\mathcal{A}}_{\beta}]=\{0\}$.
For a collection of local algebras the partial order on the indexing collection of submanifolds is containment, with $\mathcal{O}_{1}$ said to be orthogonal to $\mathcal{O}_{2}$ if the regions $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are causally independent. To obtain the corresponding notion of a von Neumann quasi-local algebra one simply replaces $\mathcal{A}$} and the ${\mathcal{A}}_{\alpha}$’s with von Neumann algebras and replaces the norm-density in (iii) with $\sigma$-weak density.
The local algebra generated by solutions to the Klein-Gordon equation and indexed by $\mathcal{K}(M,g)$ does turn out to be a quasi-local algebra. Following [3] we may construct this algebra as follows: For each $\mathcal{O}\in\mathcal{K}(M,g)$, one firstly constructs $L^{2}(\mathcal{O})$ where integration is by the volume form on $\mathcal{O}$. The algebra ${\mathcal{A}}(\mathcal{O})$ is then a copy of the $\mathrm{CCR}$ algebra constructed from Weyl operators $W(\psi)$ where the $\psi$ ranges over the solutions of the local Klein-Gordon equation on $\mathcal{O}$, namely $\{\psi\colon\psi=\widetilde{G}_{\mathcal{O}}(f),f\in\mathscr{D}(\mathcal{O},E)\}$. Here we note that the operator $\widetilde{G}_{\mathcal{O}}$ is for every $f\in\mathscr{D}(\mathcal{O},E)\}$ given by $\widetilde{G}_{\mathcal{O}}(f)=G(f_{ext})_{|\mathcal{O}}$ where $f_{ext}$ denotes the element of $\mathscr{D}(M,E)$ obtained by assigning the value 0 to all points $p\in M\backslash\mathcal{O}$. The local algebras of Brunetti, Fredenhagen and Verch [9] are very similar with the main difference being that their Weyl operators are constructed from elements of the set $\{\psi\colon\psi=\widetilde{G}_{\mathcal{O}}(f),f\in C_{0}^{\infty}(\mathcal{O})$. Readers wishing to see an axiomatic description of this approach would be well-advised to consult the papers of Fredenhagen and Rejzner [22], and Buchholz and Fredenhagen [11].
Von Neumann local algebras are typically constructed by passing to a representation by means of a regular state of the $C^{*}$-algebra setting and then taking the double commutant of the representation. At the forefront of such “good” states are the quasi-free states. Representations with these states typically lead to the iconical Araki-Woods factors. See [14] for an account of this process. However in many cases more than just quasi-freeness is needed. Hadamard states have the added advantage that they are related to Quantum Weak Energy Inequalities (QWEIs) which in turn ensure that the existence of “good” Wightman fields at the microscopic level harmonises with observance of the second law of thermodynamics at the macroscopic level. Further physical reasons also bring passive and stationary states into the picture. This topic will be explored in more detail in Part II As noted previously [14] is a good reference for information on quasi-free states and Araki-Woods factors. The paper [30] gives useful information regarding quasi-free, Hadamard and stationary states.
Besides the references mentioned above we for the sake of the reader note the following: The book of Wald [50] contains valuable information regarding relativity and differential geometry. Taking the appendices into account, the book [3] will provide the reader with a very good “quick” introduction to Lorentzian geometry and an insightful account of the Klein-Gordon and related equations. The books of Gallier and Quiantance [23, 24] give an excellent account of local flows albeit in a Riemann geometric context. Minguzzi’s notes [37] give a very comprehensive account of causality theory.
3. The standard form of von Neumann algebras
We will show that there is a categorical way to incorporate Hilbert spaces into the present framework without going to the extreme of assuming the a priori existence of some global Hilbert space of states. On a philosophical level we agree with the fact that the theory should be based on the principles of AQFT. In the modern approach to canonical quantization this means that to each quantum system there corresponds a $C^{*}$ or von Neumann algebra which encodes the information of that system. For us the basic object is however an abstract von Neumann algebra equipped with some fns weight. With this approach each such abstract von Neumann algebra describing some local region of space-time allows for the construction of a unique Hilbert space associated with that particular algebras by means of the GNS process. This Hilbert space is therefore a product of the system rather than an a priori given object.
For the reader’s convenience, we recall that the equivalent abstract definition of von Neumann algebras identifies these algebras as $C^{*}$-algebra which have a (necessarily unique) predual. In fact such abstract von Neumann algebras are the departure point for the treatment of this topic in both of [45] and [48]. What is crucial for our work is the following well known fact: any von Neumann algebra $\mathcal{M}$ (abstract or concrete) admits a faithful normal weight $\nu$ for which the associated GNS construction yields a faithful standard representation $\pi$ of $\mathcal{M}$, i.e. a normal -isomorphism into the algebra of all bounded linear operators $B(H_{\nu})$ on the GNS Hibert space $H_{\nu}$ with the image $\pi(\mathcal{M})$ being a von Neumann algebra in standard form, see either [49, Theorem IX.1.2 & Definition IX.1.13] or [25, Theorem 10.59].
If one brings the theory of the standard form of a von Neumann algebra into play, one can see that von Neumann algebras in standard form are a category which admit a tensor structure and which does not require an a priori global Hilbert space. Since the von Neumann algebra tensor product is unique there is here no selection of the “correct” tensor product that needs to be made. In addition the tensor structure associated to this category is refined enough to allow for a description of entangled states. We briefly outline these facts.
The standard form of a von Neumann algebra is an abstract characterisation of the uniqueness of the GNS-representation of the pair $({\mathcal{M}},\nu)$ where $\nu$ is a faithful normal semifinite weight on the von Neumann algebra $\mathcal{M}$}. Haagerup proved a very deep theorem essentially showing that any representation of $\mathcal{M}$} which admits objects that mimic the action of $J_{\nu}$ and $\mathscr{P}_{\nu}^{\natural}$ is a faithful copy of the GNS-representation of the pair $({\mathcal{M}},\nu).$ This claim may be made exact with the following definition:
Definition 3.1.
Given a von Neumann algebra $\mathcal{M}$} equipped with a faithful normal semifinite weight $\nu$, a quintuple $({\mathcal{M}},\pi_{0},H_{0},J,\mathscr{P})$ where $\pi_{0}$ is a faithful representation of $\mathcal{M}$} on the Hilbert space $H_{0}$, $J:H_{0}\to H_{0}$ anti-linear isometric involution, and $\mathscr{P}$ a self-dual cone of $H_{0}$, is said to be a standard form of $\mathcal{M}$} if the following conditions hold:
  • $J\pi_{0}({\mathcal{M}})J=\pi_{0}({\mathcal{M}})^{\prime}$ (the commutant of $\pi_{0}({\mathcal{M}})$),
  • $JzJ=z^{*}$ for all $z$ in the centre of $\pi_{0}({\mathcal{M}})$,
  • $J\xi=\xi$ for all $\xi\in\mathscr{P}$,
  • $a(JaJ)\mathscr{P}\subseteq\mathscr{P}$ for all $a\in\pi_{0}({\mathcal{M}})$.
(Recall that when we say that $\mathscr{P}$ is a self-dual cone, we mean that $\xi\in\mathscr{P}$ if and only if $\langle\xi,\zeta\rangle\geq 0$ for all $\zeta\in\mathscr{P}.$)
The value of the above concept is derived from the following very deep and useful theorem:
Theorem 3.2 ([27]).
The standard form of a von Neumann algebra $\mathcal{M}$} is unique in the sense that if
[…]
  • $\alpha(x)=uxu^{*}$ for $x\in\pi_{0}({\mathcal{M}})$;
  • $\widetilde{J}=uJu^{*}$;
  • $\widetilde{\mathscr{P}}=u\mathscr{P}.$
Remark 3.3.
Based on the above theorem it is clear that one may define a category VN-STDFM for which the objects are the quintuples described in Definition 3.1 and the isomorphisms unitary equivalence as described in Theorem 3.2. If both $({\mathcal{M}}_{0},H_{0},J_{0},\mathscr{P}_{0})$ and $({\mathcal{M}}_{1},H_{1},J_{1},\mathscr{P}_{1})$ are in standard form, that means that $({\mathcal{M}}_{0},H_{0},J_{0},\mathscr{P}_{0})$ must be the product of a GNS construction performed with respect to some fns weight $\nu$ on ${\mathcal{M}}_{0}$. A homomorphism from $({\mathcal{M}}_{0},H_{0},J_{0},\mathscr{P}_{0})$ to $({\mathcal{M}}_{1},H_{1},J_{1},\mathscr{P}_{1})$ would then consist of a faithful normal conditional expectation $\mathbb{E}$ satisfying $\nu\circ\mathbb{E}=\nu$ such that $({\mathcal{M}}_{1},H_{1},J_{1},\mathscr{P}_{1})$ is isomorphic to the quadruple produced by the pair $(\mathbb{E}({\mathcal{M}}_{0}),\nu_{|\mathbb{E}({\mathcal{M}}_{0})})$. The structure of these morphisms ensure that any theory constructed using this category will behave well with respect to subtheories.
3.1. VN-STDFM is a tensorial category
As can be seen from the following result, the von Neumann algebra tensor product is a tensor structure on the category VN-STDFM.
Proposition 3.4 (Prop 2a, [13]).
[…]
As can be seen from [13, Prop 2b], the above result even holds for infinite tensor products. This tensor structure even yields a prescription for combining a priori given fns weights. These facts may be deduced from [13, Prop a, b, g].
Proposition 3.5 (Prop 3g, [13]).
[…]
Let $({\mathcal{M}}\overline{\otimes}{\mathcal{N}},\pi_{{\mathcal{M}}}\otimes\pi_{{\mathcal{N}}},H_{{\mathcal{M}}}\overline{\otimes}H_{{\mathcal{N}}},J_{{\mathcal{M}}}\otimes J_{{\mathcal{N}}},\mathscr{P}^{+})$ be the standard form for ${\mathcal{M}}\overline{\otimes}{\mathcal{N}}$ described in Proposition 3.4. Then the following holds:
  • There exists a weight $\nu_{{\mathcal{M}}}\otimes\nu_{{\mathcal{N}}}$ such that $a\in\mathfrak{n}_{\nu_{{\mathcal{M}}}}$ and $b\in\mathfrak{n}_{\nu_{{\mathcal{N}}}}$ imply $a\otimes b\in\mathfrak{n}_{\nu_{{\mathcal{M}}}\otimes\nu_{{\mathcal{N}}}}$ with $\eta_{\nu_{{\mathcal{M}}}\otimes\nu_{{\mathcal{N}}}}(a\otimes b)=\eta_{\nu_{{\mathcal{M}}}}(a)\otimes\eta_{\nu_{{\mathcal{N}}}}(b)$.
  • $\Delta_{\nu_{{\mathcal{M}}}\otimes\nu_{{\mathcal{N}}}}=\Delta_{\nu_{{\mathcal{M}}}}\otimes\Delta_{\nu_{{\mathcal{N}}}}$ (equivalently $S_{\nu_{{\mathcal{M}}}\otimes\nu_{{\mathcal{N}}}}=S_{\nu_{{\mathcal{M}}}}\otimes S_{\nu_{{\mathcal{N}}}}$).
  • $(\nu_{{\mathcal{M}}}\otimes\nu_{{\mathcal{N}}})^{\prime}=(\nu_{{\mathcal{M}}})^{\prime}\otimes(\nu_{{\mathcal{N}}})^{\prime}$.
  • Let $A\subset\mathfrak{n}_{\nu_{{\mathcal{M}}}}$ and $B\subset\mathfrak{n}_{\nu_{{\mathcal{N}}}}$ be linear subspaces which are self-adjoint in the sense that $A=\{a^{*}\colon a\in A\}$ and $B=\{b^{*}\colon b\in B\}$. If $\eta_{\nu_{{\mathcal{N}}}}(A)$ and $\eta_{\nu_{{\mathcal{N}}}}(B)$ are respectively cores for $S_{\nu_{{\mathcal{M}}}}$ and $S_{\nu_{{\mathcal{N}}}}$, then $\nu_{{\mathcal{M}}}\otimes\nu_{{\mathcal{N}}}$ is the unique fns weight for which $a\otimes b\in\mathfrak{n}_{\nu_{{\mathcal{M}}}\otimes\nu_{{\mathcal{N}}}}$ for all $a\in A$ and $b\in B$.
3.2. The structure of VN-STDFM allows for entanglement
In QFT, the Reeh-Schlieder theorem implies that any chosen vector $\Psi$ can be approximated by the action of some localized operator on the vacuum. This result shows that vacuum posses small but nonvanishing and even long distance correlations which are not expected in classical Physics. So quantum correlations appeared, c.f. the Haag book [26] for details.
On the other hand AQFT, as it was mentioned, puts the emphasis on localized observables. We remind that in Quantum Theory knowledge of a system is described by a state. Thus quantum correlations should be encoded in some specific states. In particular, there are states of two (or a group) localized particles, even when the particles are separated by a large distance, which are not just a convex combination of individual quantum states. Such states are called entangled states and what is important, they describe quantum (so non-classsical) correlations. We remind that a state which is a convex combination of individual quantum states is called a separable state. It is worth pointing out that such states contain only classical correlations.
Consequently, in any axiomatic description of quantum fields, so even in Quantum Gravity, a characterization of quantum entanglement should be given. In this section we present within the framework of $W^{*}$-algebraic QFT, a general description of entangled states. This is a generalization of the characterization of entangled and separable states given in Dirac’s approach to Quantum Mechanics [32].
We begin by recalling some more facts of the theory of von Neumann algebra in the standard form. Denote by $\mathcal{M}$$_{*}$ the predual of von Neumann algebra $\mathcal{M}$}. If $\mathcal{M}$} is $\sigma$-finite, then to each $\omega\in{{\mathcal{M}}}_{*}^{+}$, there corresponds a unique $\xi\in{\mathcal{P}}$ with $\omega(\cdot)=(\xi,\cdot\xi)\equiv\omega_{\xi}(\cdot)$. (See [1] and [6], [49] for a recent account of the theory.) Furthermore, the mapping $\xi\mapsto\omega_{\xi}$ is a homeomorphism when both $\mathcal{P}$} and $\mathcal{M}$$^{+}_{*}$ are equiped with the norm topology.
[…]
Proposition 3.6 (Prop 2c, [13]).
[…]
  • If either of $\mathcal{M}$} or $\mathcal{N}$} is commutative, then $\mathscr{P}^{+}=\mathscr{P}^{+}_{0}$.
  • If both $\mathcal{M}$} and $\mathcal{N}$} are noncommutative, then $\mathscr{P}^{+}\neq\mathscr{P}^{+}_{0}$.
The striking conclusion for physics is: separable normal states of composition system, describing classical correlations, form only a subset in the set of normal states. Consequently, there are states $\mathscr{P}^{+}\setminus\mathscr{P}^{+}_{0}$, entangled states, which encode non-classical - quantum correlations. Keeping the Reeh-Schlieder theorem in mind, this result is not surprising. A system composed of two quantum subsystems must have quantum correlations. However, the great advantage of the above result is the description of states with quantum correlations.
It is worth pointing out that the above nice result was obtained in the framework of $W^{*}$-algebras. This is huge advantage over the $C^{*}$-approach to QFT.
Part I Quantum Killing vector fields for AQFT
In the brief comment near the end of the introduction regarding $L^{\cosh-1}$-regularity we already indicated the utility of quantum integration theory for QFT, albeit for the special relativistic setting. We shall revisit this issue for general spacetimes in the next part. For now we pass to exploring differential structures for curved spacetimes. In so doing we shall in this part primarily focus on the $\mathrm{CCR}$ local algebras described in [3]. To be more specific about differentiation, more specific algebras and states need to be considered. For now we content ourselves with sketching the broad panorama that begs further analysis.
4. Killing local flows on compacta of Lorentzian manifolds
The starting point of our analysis is essentially a localisation of Proposition 2.3(b) of [9].
Definition 4.1.
Let $(M,g)$ be a time-oriented Lorentzian manifold. A tangent vector $v$ is causal if $v\neq 0$ and $g(v,v)\leq 0$. A curve is called causal if all of its tangent vectors are causal. The causal curves between two points in a set $V$ is usually denoted by $J^{+}(V)\cap J^{-}(V)$ where $J^{+}(V)=\cup_{p\in V}J^{+}(p)$ and $J^{+}(p)$ is the causal future (resp. $J^{-}(p)$ the causal past) of the point $p$, etc.
A subset $V\subset M$ is said to be causally convex if it contains all causal curves between two points in the set, that is if $J^{+}(V)\cap J^{-}(V)\subset V$. For an arbitrary subset $V\subset M$ the causally convex hull of $V$ is defined to be the union of the set of the images of all causal curves which start and end in $V$. We shall however write $\mathrm{cau}(V)$ for the causal hull.
A subset $\Omega$ of $(M,g)$ is called causally compatible if for all points $x\in\Omega$ we have $J^{\pm}_{\Omega}(x)=J_{M}^{\pm}(x)\cap\Omega$.
Note that the inclusion $J^{\pm}_{\Omega}(x)\subseteq J_{M}^{\pm}(x)\cap\Omega$ always holds. The condition of being causally compatible therefore means that whenever two points in $\Omega$ can be joined by a causal curve in $M$, this can also be done inside $\Omega$.
Notation: We will denote the class of all relatively compact causally compatible globally hyperbolic open subsets of a globally hyperbolic spacetime $(M,g)$ by $\mathcal{I}_{0}(M,g)$. Bär, Ginoux and Pfäffle [3] use the notation $\mathcal{I}(M,g)$ for $\mathcal{I}_{0}(M,g)\cup\{\emptyset,M\}$ added. Both these classes are clearly subclasses of the class of all open relatively compact causally convex subsets of $M$. We follow [9] by denoting the set of all open relatively compact causally convex subsets of $M$ by $\mathcal{K}(M,g)$. Elements of this class can all be shown to be globally hyperbolic in their own right
It is of independent interest to note that global hyperbolicity may be characterised in terms of compactness of the causally convex hull,
Proposition 4.2.
The following are equivalent for a Lorentzian manifold:
  • The manifold is globally hyperbolic.
  • The manifold is causal and the causally convex hull of any compact subset is compact. [28, Proposition 2.5]
Both imply that the causally convex hull of any relatively compact set is relatively compact. [37, Page 4]
We are particularly concerned with Killing vector fields on Lorentzian manifolds. These are (smooth) vector fields $X$ for which the Lie derivative of the metric tensor vanishes, that is $L_{X}g=0$. The primary reason for our interest stems from the following fact
Remark 4.3.
It is well-known that the Killing vector fields on Minkowski space correspond exactly to the action of the Poincaré group on Minkowski space. So for Lorentzian manifolds these fields may be viewed as a curved analogue of the action of the Poincaré group on Minkowski space.
Killing vector fields are very beautifully characterised as those vector fields with isometric local flows. For the sake of the reader we record the essentials.
Definition 4.4.
Let $M$ and $N$ be Lorentzian manifolds with metric tensors $g_{1}$ and $g_{2}$. An isometry from $M$ to $N$ is a diffeomorphism $\Psi:M\to N$ that preserves the metric tensor, that is $\Psi^{*}(g_{2})=g_{1}$. Explicitly, $\langle d\Psi_{p}(v),d\Psi_{p}(w)\rangle_{\Psi}(p)=\langle v,w\rangle_{p}$ for all $v,w\in T_{p}(M)$. For Lorentzian manifolds this means that for each $p$ in the domain of $\Psi$, the map $d\Psi_{p}$ is a Lorentzian isometry from $T_{p}(M)$ to $T_{\Psi(p)}(M)$. (See the comment after Definition 3.6 of [40].)
Theorem 4.5.
A vector field $Z$ on a semi-Riemannian manifold is a Killing vector field iff the stages $\Psi_{s}^{Z}$ of the local flow are isometries. ([40, 9.21])
Note: Lemma 4.4.8 of [3] is very important. It shows that for globally hyperbolic spacetimes the relatively compact causally compatible globally hyperbolic open subsets is a directed set with an orthogonality relation. However more is true. What is actually proved is the following:
Proposition 4.6.
Let $(M,g)$ be a globally hyperbolic Lorentzian manifold. For any compact subset $K$ there exists some $\mathcal{O}\in\mathcal{K}(M,g)$ containing $K$. Moreover the class of subsets $\mathcal{K}(M,g)$ is a directed set with orthogonality relation. (Compare [3, Lemma A.5.11].) Finally for any $\mathcal{O}_{1}\in\mathcal{I}_{0}(M,g)$ there exists $\mathcal{O}_{2}\in\mathcal{K}(M,g)$ such that $\mathcal{O}_{1}\subset\mathcal{O}_{2}$.
Proof.
The proof of [3, Proposition A.5.13] relies on Proposition A.5.12 of [3]. However the hypothesis of [3, Proposition A.5.12] understates what is actually proven. In the second sentence of that proof the authors actually show that the set constructed there is not just causally compatible, but in fact causally convex. So the set constructed in [3, Proposition A.5.13] must also be causally convex since it is constructed by an application of [3, Proposition A.5.12]. The first claim therefore follows from [3, Proposition A.5.13]. The second claim follows from [3, Lemma 4.4.8]. To see this note that the proof of [3, Lemma 4.4.8] relies on an application of [3, A.5.11, A.5.12 & A.5.13]. We have already seen that [3, A.5.12 & A.5.13] actually yield elements of $\mathcal{K}(M,g)$. A consideration of the proof of [3, Lemma A.5.11] shows that here too the authors actually prove causal convexity and not just causal compatibility. Hence the claim follows. The final claim follows by applying the first assertion to the fact that $\overline{\mathcal{O}}_{1}$ is compact. ∎
Corollary 4.7.
Let ${\mathcal{A}}(M)$ be a local algebra of a globally hyperbolic spacetime $(M,g)$ defined as in Lemma 4.4.8 and Corollary 4.4.12 of [3]. Then $\cup_{\mathcal{K}(M,g)}{\mathcal{A}}(\mathcal{O})=\cup_{\mathcal{I}_{0}(M,g)}{\mathcal{A}}(\mathcal{O})$.
Proof.
This is a straightforward consequence of the final claim of the proposition and the fact that $\mathcal{K}(M,g)\subset\mathcal{I}_{0}(M,g)$. ∎
Corollary 4.8.
Let $Z$ be an arbitrary smooth vector field and $K$ a compact neighbourhood. Then there exists some $t_{K}>0$ such that $K$ is in the domain of each $\Psi_{s}$ with $0<|s|\leq t_{K}$. Given any relatively compact causally compatible globally hyperbolic open subset $\mathcal{O}$ there exists some $t_{\mathcal{O}}>0$ such that $\mathcal{O}$ is in the domain of each $\Psi_{s}$ with $0<|s|\leq t_{\mathcal{O}}$.
Proof.
The first part follows from a compactness argument. First consider positive indices $t_{p}$. Each $p\in K$ is an element of $\mathrm{dom}(\Psi_{t_{p}}^{Z})$ for $t_{p}>0$ small enough. By compactness $K$ can be covered by finitely many of the $\mathrm{dom}(\Psi_{t_{p}}^{Z})$’s. Since the domains increase as $t$ decreases to 0, taking the minimum of the remaining $t$’s yields a $t>0$ for which $K\subset\mathrm{dom}(\Psi_{s}^{Z})$ for each $0<s<t$. Now repeat the argument for negative $t$’s. For the second claim simply apply the above to the closure of $\mathcal{O}$. ∎
Lemma 4.9.
Let $(M,g)$ be a time-oriented Lorentzian manifold and $Z$ a Killing vector field with $\Psi_{s}^{Z}$ the stages of the local flow. Then $\Psi_{s}^{Z}$ will map causal curves between two points in its domain to causal curves.
Proof.
Let $g$ be the canonical Lorentzian-Riemannian metric tensor on $M$ and $Z$ a Killing field with stages $\Psi_{s}^{Z}$. For part 1 all but the claim about preservation of causal compatibility are a consequence of continuity. Given a causally compatible subset $K$ of the domain of $\Psi_{s}^{Z}$, the claim about causal compatibility will follow if we can show that $\Psi_{s}^{Z}$ (and therefore also $\Psi_{-s}^{Z}=(\Psi_{s}^{Z})^{-1}$) maps causal curves to causal curves. Let $t\to x(t)$ $(a\leq t\leq b)$ be a causal curve in $K$. Since by definition the stages $\Psi_{s}^{Z}$ of the local flow of $Z$ are isometries we surely have that $0\geq\langle x^{\prime}(t_{0}),x^{\prime}(t_{0})\rangle_{x(t_{0})}=\lim_{t\to 0}\frac{1}{t^{2}}\langle x(t_{0}+t)-x(t_{0}),x(t_{0}+t)-x(t_{0})\rangle_{x(t_{0})}=\lim_{t\to 0}\frac{1}{t^{2}}\langle d(\Psi_{s})_{x(t_{0})}(x(t_{0}+t))-d(\Psi_{s})_{x(t_{0})}(x(t_{0})),d(\Psi_{s})_{x(t_{0})}(x(t_{0}+t))-d(\Psi_{s})_{x(t_{0})}(x(t_{0}))\rangle_{\Psi_{s}(x(t_{0}))}=\langle d(\Psi_{s})_{x(t_{0})}(x^{\prime}(t_{0})),d(\Psi_{s})_{x(t_{0})}(x^{\prime}(t_{0}))\rangle_{\Psi_{s}(x(t_{0}))}$. So stages of local flows of Killing fields preserve causal curves. ∎
Proposition 4.10.
Let $\Psi_{s}^{Z}$ be the stages of a local flow of a Killing vector field $Z$. For any $\mathcal{O}\in\mathcal{K}(M,g)$ there exists some $0<t_{0}$ so that $\Psi_{s}^{Z}$ is defined on $\mathcal{O}_{\cup}=\cup_{s\in[-t_{0},t_{0}]}\Psi_{s}^{Z}(\mathcal{O})$ for all $s\in[-t_{0},t_{0}]$ with $\Psi_{s+r}^{Z}=\Psi_{s}^{Z}\Psi_{r}^{Z}$ on $\mathcal{O}$ whenever $t_{0}<s,r,s+r<t_{0}$ and with each $\Psi_{s}$ $(-t_{0}<s<t_{0})$ mapping $\mathcal{O}$ onto another element of $\mathcal{K}(M,g)$.
Proof.
The most difficult property to prove is the preservation of causal convexity for $t$ in some interval. The rest of the claims are fairly direct consequences of the description in for example §9.3 of [23]. We proceed with proving the claim regarding causal convexity. Let $K$ be a compact set and select $t>0$ so that $\Psi_{s}^{Z}$ is defined on $K$ for all $s\in[-t,t]$ with $\Psi_{s+r}^{Z}=\Psi_{s}^{Z}\Psi_{r}^{Z}$ on $K$ whenever t<s,r,s+r<t-t<s,r,s+r<t. By Tychonoff’s theorem $\Pi_{s\in[-t,t]}\Psi_{s}^{Z}(K)$ is compact. We recall that $\Pi_{s\in[-t,t]}\Psi_{s}^{Z}(K)$ is equipped with the coarsest topology for which all the projections $\pi_{t}\colon\Pi_{s\in[-t,t]}\Psi_{s}^{Z}(K)\to\Psi_{t}^{Z}(K)$ are continuous [5, Example 3, Ch 1, §2.3]. That in turn ensures that the function $\iota\colon\Pi_{s\in[-t,t]}\Psi_{s}^{Z}(K)\to\cup_{s\in[-t,t]}\Psi_{s}^{Z}(K)$ which for every $x\in\Psi_{t}^{Z}(K)$ is given by $\iota(t,x)=x\in\Psi_{t}^{Z}(K)$, is a continuous surjection. But then the image $\cup_{s\in[-t,t]}\Psi_{s}^{Z}(K)$ is itself compact. We may take $K$ to be the closure of some $\mathcal{O}\in\mathcal{K}(M,g)$. By Proposition 4.6 we may then select another element $\mathcal{O}_{\cup}\in\mathcal{K}(M,g)$ containing $\cup_{s\in[-t,t]}\Psi_{s}^{Z}(K)$. Now pick $0<t_{0}<t$ so that $\Psi_{s}^{Z}$ is defined on $\mathcal{O}_{\cup}$ for all $s\in[-t_{0},t_{0}]$ with $\Psi_{s+r}^{Z}=\Psi_{s}^{Z}\Psi_{r}^{Z}$ on $\mathcal{O}_{\cup}$ whenever $t_{0}<s,r,s+r<t_{0}$. Given $p,q\in\mathcal{O}$ the causal convexity of $\mathcal{O}_{\cup}$ ensures that any causal curve $\gamma$ joining $\Psi_{t_{0}}^{Z}(p)$ and $\Psi_{t_{0}}^{Z}(q)$ is contained in $\mathcal{O}_{\cup}$. Since $\Psi_{-t_{0}}^{Z}$ is defined on $\mathcal{O}_{\cup}$ it will map the causal curve $\gamma$ onto a causal curve $\Psi_{-t_{0}}^{Z}(\gamma)$ joining $p$ and $q$. The causal convexity of $\mathcal{O}$ ensures that this curve is contained in $\mathcal{O}$ joining $p$ and $q$. But then $\gamma=\Psi_{t_{0}}^{Z}(\Psi_{-t_{0}}^{Z}(\gamma))$ is a causal curve inside $\Psi_{t_{0}}^{Z}(\mathcal{O})$ joining $\Psi_{t_{0}}^{Z}(p)$ and $\Psi_{t_{0}}^{Z}(q)$. So $\Psi_{t_{0}}^{Z}(\mathcal{O})$ is causally convex. ∎
5. Killing flows on local algebras: an illustrative example based on the Klein-Gordon equation
Remark 5.1.
We will use the ideas in Lemma 4.4.8 and Corollary 4.4.12 of [3] to construct a local algebra from the collection $\mathcal{K}(M,g)$. Each local algebra ${\mathcal{A}}(\mathcal{O})$ will be the Weyl algebra corresponding to the solutions of the localised version of the Klein-Gordon equation as described in Proposition 3.5.1 of [3]. We remind the reader that a Weyl algebra is a unital $C^{*}$-algebra generated by a set of unitaries $\{W(f)\colon f\in{\mathcal{H}}\}$ governed by the relations $W(f)W(g)=e^{-i\sigma(f,g)/2}W(f+g)$ and $W(f)^{*}=W(-f)$, where $\mathcal{H}$} is a real symplectic space equipped with a non-singular symplectic form $\sigma$.
To see the claim that the local algebras mentioned above are Weyl algebras of solutions of the Klein-Gordon equation, note that Theorem 3.2.15 of [2] confirms that in the globally hyperbolic case the quotient space described prior to [3, Lemma 4.3.8] may be thought of as a space of solutions of the Klein-Gordon equation. Let $\mathcal{O}$, $t_{0}$ and $\Psi_{s}^{Z}$ be as in the previous Proposition. The remaining challenge is to show that for any stage $\Psi_{s}^{Z}$ $(0<s\leq t_{0})$, we will for the local algebras constructed according to the prescription in Lemma 4.4.8 and Corollary 4.4.12 of [3] have that the map $\Psi_{s}^{Z}\colon\mathcal{O}\to M$ canonically lifts to a *-isomorphism $\pi_{s}^{Z}$ from ${\mathcal{A}}(\mathcal{O})$ to ${\mathcal{A}}(\Psi_{s}^{Z}(\mathcal{O}))$ with $\pi_{s}^{Z}\circ\pi_{r}^{Z}=\pi_{r+s}^{Z}$ whenever the same is true for the $\Psi_{s}^{Z}$’s. This seems to only be possible for the class of sets $\mathcal{K}(M,g)$ described in [9].
Note: For the rest of this section ${\mathcal{A}}(M)$ will be the local algebra constructed from the collection $\mathcal{K}(M,g)$ according to the ideas in Lemma 4.4.8 and Corollary 4.4.12 of [3]
Lemma 5.2.
Let $(M,g)$ be globally hyperbolic and let $\Psi$ be an isometric diffeomorphism which maps $\mathcal{O}\in\mathcal{K}(M,g)$ onto another element $\Psi(\mathcal{O})$ of $\mathcal{K}(M,g)$. Let $P_{\mathcal{O}}$ and $P_{\Psi(\mathcal{O})}$ be the differential operators corresponding to the Klein-Gordon equation on $\mathcal{O}$ and $\Psi(\mathcal{O})$ respectively. Let $\widetilde{G}_{\mathcal{O}}^{\pm}$ (resp. $\widetilde{G}_{\Psi(\mathcal{O})}^{\pm}$) be the corresponding Green’s operators as described in [3, Prop 3.5.1].
  • If $\Psi^{-1}$ preserves time-orientation, then $f\to\widetilde{G}^{+}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi$ (where $f\in\mathscr{D}(\mathcal{O},E)$) agrees with $\widetilde{G}^{+}_{\mathcal{O}}$.
  • If $\Psi^{-1}$ reverses time-orientation, then $f\to\widetilde{G}^{+}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi$ (where $f\in\mathscr{D}(\mathcal{O},E)$) agrees with $\widetilde{G}^{-}_{\mathcal{O}}$.
Proof.
We observe that $\mathcal{O},\Psi(\mathcal{O})\in\mathcal{K}(M,g)$ are in effect globally hyperbolic Lorentzian manifolds in their own right. We therefore have access to the results of Dimock [15] who showed on page 226 of his paper that $\Psi^{*}P_{\mathcal{O}}=P_{\Psi(\mathcal{O})}\Psi^{*}$ where $\Psi^{*}$ is the push-forward. Given $f\in\mathscr{D}(\mathcal{O},E)$ this will by [3, Def 3.4.1] mean that
$\widetilde{P}_{\mathcal{O}}(\widetilde{G}^{\pm}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi)=P_{\Psi(\mathcal{O})}\widetilde{G}^{\pm}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi$
$PG(f\circ\Psi^{-1}_{ext})_{|\Psi(\mathcal{O})}\circ\Psi=(f\circ\Psi^{-1}_{ext})_{|\Psi(\mathcal{O})}\circ\Psi=f.$
Similarly
$\widetilde{G}^{\pm}_{\Psi(\mathcal{O})}(\widetilde{P}_{\mathcal{O}}(f)\circ\Psi^{-1})\circ\Psi=\widetilde{G}^{\pm}_{\Psi(\mathcal{O})}(\widetilde{P}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}))\circ\Psi$
$G(P(f\circ\Psi^{-1})_{ext})\circ\Psi=(f\circ\Psi^{-1})_{ext}\circ\Psi=f.$
If $\Psi^{-1}$ preserves time-orientation, then since it also preserves causal curves, we may therefore conclude from [3, Prop 3.5.1] that
$\mathrm{supp}(\widetilde{G}^{\pm}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi)\subseteq\Psi^{-1}(\mathrm{supp}(\widetilde{G}^{\pm}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})))\subset\Psi^{-1}(J^{\pm}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}))=J^{\pm}_{\mathcal{O}}(f).$
Thus $f\to\widetilde{G}^{\pm}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi$ is an advanced Green’s operator for the action of $P_{\mathcal{O}}$ on $\mathcal{O}$. But for globally hyperbolic Lorentzian manifolds such operators are unique, which then proves the claim in this case. The proof of the case where $\Psi^{-1}$ reverses time-orientation is similar. ∎
Theorem 5.3.
Let $(M,g)$ be a globally hyperbolic and let $\Psi$ be an isometric diffeomorphism which maps $\mathcal{O}\in\mathcal{K}(M,g)$ onto another element $\Psi(\mathcal{O})$ of $\mathcal{K}(M,g)$.
  1. (1)
    For the local algebra described in part (1) of [31, Theorem 3.4] the prescription $f\to f\circ\Psi^{-1}$ where $f\in C^{\infty}_{0}(\mathcal{O})$ lifts to a *-isomorphism $\pi_{\Psi}$ from ${\mathcal{A}}(\mathcal{O})$ to ${\mathcal{A}}(\Psi(\mathcal{O}))$ which sends $W(f)$ where $f\in C^{\infty}_{0}(\mathcal{O})$ to $W(f\circ\Psi^{-1})$.
  2. (2)
    If $\Psi^{-1}$ preserves time-orientation then for the local algebra described in [3, Lemmata 4.4.8 & 4.4.10] we also have that the map $\Psi\colon\mathcal{O}\to\Psi(\mathcal{O})$ canonically lifts to a *-isomorphism $\pi_{\Psi}$ from ${\mathcal{A}}(\mathcal{O})$ to ${\mathcal{A}}(\Psi(\mathcal{O}))$ which implements $\Psi$ in the sense of sending $W(\widetilde{G}_{\mathcal{O}}(f))$ where $f\in\mathscr{D}(\mathcal{O},E)$ to $W(\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}))$.
  3. (3)
    If $\Psi^{-1}$ reverses time-orientation then for the local algebra described in [3, Lemmata 4.4.8 & 4.4.10] we have that the map $\Psi\colon\mathcal{O}\to\Psi(\mathcal{O})$ canonically lifts to a *-anti-isomorphism $\pi_{\Psi}$ from ${\mathcal{A}}(\mathcal{O})$ to ${\mathcal{A}}(\Psi(\mathcal{O}))$ which implements $\Psi$ in the above sense, but this time sending $W(\widetilde{G}_{\mathcal{O}}(f))$ to $W(-\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}))$.
Proof.
Claim 1: Firstly note that [24, Prop 5.19] suggests that if $d\mu$ is the measure coming from the volume form, then on $\mathcal{O}$ the stage $\Psi_{s}$ is measure preserving. So for any $f,g\in C_{0}^{\infty}(M)$ with support in $\mathcal{O}$ we have that $\int_{\mathcal{O}}f\overline{g}\,d\mu=\int_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})(\overline{g\circ\Psi^{-1}})\,d\mu$ where we have used the fact that $f\circ\Psi^{-1}$ and $g\circ\Psi^{-1}$ are elements of $C_{0}^{\infty}(M)$ supported on $\Psi(\mathcal{O})$. The claim now follows from [3, Theorem 4.2.9].
Claims 2 and 3: Suppose that $\Psi$ either reverses or preserves time-orientation, and let $\widetilde{P}_{\mathcal{O}}$ be the Klein-Gordon operator and $\widetilde{G}_{\mathcal{O}}$ the Green’s operator associated with $\mathcal{O}$ as described in [3, Prop 3.5.1]. It is now clear from Lemma 5.2 that the prescription $f\to f\circ\Psi^{-1}$ will map $\widetilde{G}^{+}_{\mathcal{O}}(f)-\widetilde{G}^{-}_{\mathcal{O}}(f)$ onto either $\widetilde{G}^{+}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi-\widetilde{G}^{-}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi$ or $\widetilde{G}^{-}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi-\widetilde{G}^{+}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi$, depending on whether $\Psi$ either preserves or reverses time-orientation.
Case 1 - preservation of time-orientation: If $\Psi$ preserves time orientation then the facts about preservation of measure noted in the first part of the proof, ensure that
$\int_{M}\langle g,\widetilde{G}_{\mathcal{O}}(f)\rangle\,dV$ $=$ $\int_{M}\langle g,\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\circ\Psi\rangle\,dV$
$=$ $\int_{M}\langle g\circ\Psi^{-1},\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\rangle\,dV.$
Since $\Psi$ preserves time-orientation, the prescription $\psi\to\psi\circ\Psi^{-1}$ yields a symplectic map mapping $\{\widetilde{G}_{\mathcal{O}}(f)\colon f\in\mathscr{D}(\mathcal{O},E)\}$ onto $\{\widetilde{G}_{\Psi(\mathcal{O})}(f)\colon f\in\mathscr{D}(\Psi(\mathcal{O}),E)\}$. To in this case obtain the conclusion, all that is required is to apply Theorem 4.2.9 of [3] to the above fact.
Case 2 - reversal of time-orientation: If on the other hand $\Psi$ reverses time-orientation, a similar computation shows that $\int_{M}\langle g,\widetilde{G}_{\mathcal{O}}(f)\rangle\,dV=-\int_{M}\langle g\circ\Psi^{-1},\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\rangle\,dV$. To obtain the conclusion we will in this case need to work a bit harder. So here the prescription $f\to f\circ\Psi^{-1}$ yields a *-isomorphism from ${\mathcal{A}}(\mathcal{O})$ to the algebra $\widetilde{{\mathcal{A}}}(\Psi(\mathcal{O}))$ constructed from the degenerate symplectic form $(\varphi,\varrho)\to-\int_{M}\langle\varphi,\widetilde{G}_{\Psi(\mathcal{O})}(\varrho)\rangle\,dV$ on $\mathscr{D}(\Psi(\mathcal{O}),E)$, which maps $W(\widetilde{G}_{\mathcal{O}}(f))$ to $\widetilde{W}(\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}))$ (where $\widetilde{W}(\varphi)$ are the Weyl operators generated using the above symplectic form. A careful consideration of the third displayed equation in [3, Example 4.2.2] shows that for any $\varphi\in\mathscr{D}(\Psi(\mathcal{O}),E)$, the Weyl operator $\widetilde{W}(\varphi)$ described above corresponds to $W(-\varphi)=W(\varphi)^{*}$. Thus the algebras ${\mathcal{A}}(\Psi(\mathcal{O}))$ and $\widetilde{{\mathcal{A}}}(\Psi(\mathcal{O}))$ agree as they are generated by the same operators. The *-isomorphism obtained above then sends $W(\widetilde{G}_{\mathcal{O}}(f))$ to $W(-\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}))=W(\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}))^{*}$. By now following this *-isomorphism with involution, we obtain the promised *-anti-isomorphism. (Some technicalities have been suppressed for the sake of clarity. For full details see pages 129-133 of [3].) ∎
Remark 5.4.
If $(M,g)$ is connected then all local diffeomorphisms either preserve or reverse time-orientation. See the discussion preceding [40, Cor 7.9].
Lemma 5.5.
Let $\Psi_{s}^{Z}$ be the stages of a local flow of a Killing vector field $Z$ and let $\mathcal{O}\in\mathcal{K}(M,g)$ be given. Let $0<t_{0}$ be as in Propositions 4.10. For any $s\in[-t_{0},t_{0}]$ the stage $\Psi_{s}^{Z}$ will then preserve time-orientation on $\mathcal{O}$.
Proof.
We remind the reader that $\mathcal{O}$ is globally hyperbolic in its own right and that globally hyperbolic Lorentzian manifolds are connected by definition (see [3]). It is clear from the preceding remark that $\Psi_{s/2}^{Z}$ will for any $s\in[-t_{0},t_{0}]$ either preserve or reverse time-orientation. But in either case $\Psi_{s}^{Z}=\Psi_{s/2}^{Z}\circ\Psi_{s/2}^{Z}$ will then preserve time-orientation. The claim therefore follows. ∎
Theorem 5.3 and Lemma 5.5 now yields the following result:
Corollary 5.6.
Let $(M,g)$ be a globally hyperbolic manifold and let $Z$ be a Killing vector field. Let $\mathcal{O}\in\mathcal{K}(M,g)$. Select a positive real $t_{0}>0$ so that $\Psi_{s}^{Z}$ is defined on $\mathcal{O}_{\cup}$ for all $s\in[-t_{0},t_{0}]$ with $\Psi_{s+r}^{Z}=\Psi_{s}^{Z}\Psi_{r}^{Z}$ on $\mathcal{O}_{\cup}$ whenever $t_{0}<s,r,s+r<t_{0}$ and with each $\Psi_{s}^{Z}$ $(-t_{0}<s<t_{0})$ mapping $\mathcal{O}$ onto a relatively compact causally compatible globally hyperbolic open subset of $M$.
  1. (1)
    For any stage $\Psi_{s}^{Z}$ $(0<s\leq t_{0})$ there exists a * isomorphism $\pi_{s}^{Z}$ from ${\mathcal{A}}(\mathcal{O})$ to ${\mathcal{A}}(\Psi_{s}^{Z}(\mathcal{O}))$ which implements the action of the local flow at the algebraic level.
  2. (2)
    On passing to the Fock representation of these Weyl algebras, we note that for any $a\in{\mathcal{A}}(\mathcal{O})$ we have that $\pi_{s}(Z)(a)$ strongly converges to $a$ as $s\to 0$.
Proof.
Only the final claim needs to be proven. We know from [3, Proposition 4.6.10] that $\pi_{s}^{Z}(W(\widetilde{G}_{\mathcal{O}}(f)))=W(\widetilde{G}_{\Psi_{s}^{Z}(\mathcal{O})}(f\circ(\Psi_{s}^{Z})^{-1}))=W(\widetilde{G}_{\mathcal{O}}(f)(\Psi_{s}^{Z})^{-1})$ is strongly convergent to $W(\widetilde{G}_{\mathcal{O}}(f))$ on Fock space as $s\to 0$. The same is therefore true for the span of the Weyl operators, which is norm dense in the $\mathrm{CCR}$ algebra ${\mathcal{A}}(\mathcal{O})$. So by suitably approximating we can show that for any $a\in{\mathcal{A}}(\mathcal{O})$, $\pi_{s}^{Z}(a)\to a$ strongly as $s\to 0$. ∎
6. Quantum Killing Lie derivatives
We shall here pass to the von Neumann algebra setting assuming all our von Neumann local algebras ${\mathcal{M}}(\mathcal{O})$ to be the $\sigma$-weak closure of the matching $C^{*}$-algebras ${\mathcal{A}}(\mathcal{O})$ in some appropriate representation. (We shall elaborate more on what we mean by “appropriate representations” in the next part of the paper.) In the notation of Corollary 5.6 we shall also assume that for any $a\in{\mathcal{M}}(\mathcal{O})$ we have that $\pi_{s}^{Z}(a)$ strongly converges to $a$ as $s\to 0$. To see that this holds in the Fock space case, notice that by [14, Theorem 6] the Bogoliubov transformations which are used to construct the $\pi^{Z}_{t}$s are implemented by a unitary. Thus the $\pi_{t}^{Z}$s admit a canonical extension to the double commutants of the ${\mathcal{A}}(\mathcal{O})$s. In addition since each of the $\pi_{t}^{Z}$s described in Corollary 5.6 is contractive, the convergence of $(\pi_{t}^{Z}(a))$ noted there is actually in the $\sigma$-strong topology. By convexity each ${\mathcal{A}}(\mathcal{O})$ is $\sigma$-strongly dense in ${\mathcal{M}}(\mathcal{O})$, and hence in this case the convergence we are assuming is just a continuous extension of the convergence noted in part (2) of Corollary 5.6. We pass to considering the consequences of strong continuity at 0. Formally Quantum Killing Lie derivatives are the infinitesimal generators of the quantum local flows of Killing vector fields. These quantum Lie derivatives are defined as follows:
Definition 6.1.
Let $(M,g)$ be globally hyperbolic. Given a Killing vector field $Z$ on $(M,g)$, we define $\delta_{Z}$ to be the operator $\cup_{\mathcal{O}\in\mathcal{K}(M,g)}{\mathcal{M}}(\mathcal{O})\to{\mathcal{M}}(M)$ with domain $\mathrm{dom}(\delta_{Z})=\{f\in{\mathcal{M}}(\mathcal{O})\colon\mathcal{O}\in\mathcal{K}(M,g),\lim_{t\to 0}\frac{1}{t}(\pi_{t}^{Z}(f)-f)\mbox{ exists as an element of }{\mathcal{M}}(M)\}$ and with values given by $\delta_{Z}(f)=\lim_{t\to 0}\frac{1}{t}(\pi_{t}^{Z}(f)-f)$.
Some technical background is required for us to get positive results.
6.1. Generators of one-parameter local groups of contractions
We are ultimately interested in describing the generators of the quantum Killing flows described in the previous subsection. To be able to achieve that objective the appropriate machinery needs to be developed. This subsection is devoted to laying that groundwork.
Definition 6.2.
We call a set $\{T_{t}\colon t\in\mathbb{R}\}$ of partially defined linear contractions on a von Neumann algebra $\mathcal{M}$} a strongly continuous one-parameter local group of contractions if there exists an associated collection $\{D_{\alpha}\colon\alpha\in(0,\infty)\}$ of subspaces such that
  1. (1)
    $D_{\alpha}\subseteq D_{\beta}$ if $0<\beta\leq\alpha$ with $\cup_{\alpha>0}D_{\alpha}$ strongly dense in $\mathcal{M}$};
  2. (2)
    $D_{\alpha}\subseteq\mathrm{dom}(T_{t})$ if $|t|\leq\alpha$ with $T_{0}=\mathrm{Id}$;
  3. (3)
    for every $x\in D_{\alpha}$ we have $T_{s}T_{t}(x)=T_{s+t}(x)$ whenever $s,t,s+t$ all belong to $\alpha,\alpha]$;
  4. (4)
    for every $x\in\cup_{\alpha>0}D_{\alpha}$ we have that $T_{t}(x)\to x$ strongly as $t\to 0$.
Note that for large $\alpha>0$, the subspace $D_{\alpha}$ may just be $\{0\}$.
Remark 6.3.
Since for each $x\in D_{\alpha}$ the set $\{T_{t}(x)\colon t\in(-\alpha,\alpha)\}$ is bounded, the convergence in part (4) above is actually $\sigma$-strong. If the $T_{t}$s preserve adjoints the convergence is even $\sigma$-strong*.
Lemma 6.4.
Let $\{T_{t}\colon t\in\mathbb{R}\}$ be a strongly continuous 1-parameter local group of contractions on $\mathcal{M}$}. For every $\alpha>0$ and every $x\in D_{\alpha}$ the function $t\to T_{t}(x)$ is then a point to $\sigma$-strong continuous function from $\alpha,\alpha]$ to $\mathcal{M}$}.
Proof.
A straightforward modification of the proof of [41, Corollary 2.3] suffices to prove the result. Continuity at $t=0$ follows from the definition. For $t\in(-\alpha,\alpha)$ we may simply replace $x$ with $T_{t}(x)$ and apply the claims in the definition to the limit $\lim_{s\to 0}T_{s}(T_{t}(x))$. The endpoints need to be handled with more care as one needs to check for one-sided continuity. If for example $t=-\alpha$ then for $s>0$ small enough we will for any $\xi\in H$ have that $\|T_{-\alpha}(x)\xi-T_{-\alpha+s}(x)\xi\|\leq\|T_{-\alpha}\|.\|x(\xi)-T_{s}(x)(\xi)\|\to 0$ as $s\searrow 0$. ∎
Lemma 6.5.
Let $\{T_{t}\colon t\in\mathbb{R}\}$ be a strongly continuous 1-parameter local group of contractions on $\mathcal{M}$} and let $A$ be its infinitesimal generator. For every $x\in\cup_{\alpha>0}D_{\alpha}$ the following will hold for every $t$ with $|t|$ small enough:
  1. (1)
    $\lim_{h\searrow 0}\frac{1}{h}\int_{t}^{t+h}T_{s}(x)\,ds=\lim_{h\nearrow 0}\int_{t+h}^{t}T_{s}(x)\,ds=T_{t}(x)$.
  2. (2)
    $\int_{0}^{t}T_{s}(x)\,ds\in\mathrm{dom}(A)$ with $A\int_{0}^{t}T_{s}(x)\,ds=T_{t}(x)-x$.
Proof.
Given $x\in\cup_{\alpha>0}D_{\alpha}$ we clearly have that $x\in D_{\beta}$ for some $\beta>0$. In view of the continuity noted in Lemma 6.4, part (1) follows from a straightforward generalisation a well known fact regarding the Riemann integral . It remains to prove (2). Given $t\in(0,\beta)$, it then follows from Definition 6.2 and Lemma 6.4 that we will for any $h\in(0,\beta-t)$ which is small enough so that $|h|<|t|$ have that
$\frac{1}{h}(T_{h}-{\mathds{1}})\int_{0}^{t}T_{s}(x)\,ds=\frac{1}{h}\int_{0}^{t}(T_{s+h}-T_{s})(x)\,ds=\frac{1}{h}\left[\int_{t}^{t+h}T_{s}(x)\,ds-\int_{0}^{h}T_{s}(x)\,ds\right].$
As $h\searrow 0$, the right hand side will by part (1) converge to $T_{t}(x)-x$. The claim now holds by definition. ∎
Corollary 6.6.
Let $\{T_{t}\colon t\in\mathbb{R}\}$ be a strongly continuous 1-parameter local group of contractions on $\mathcal{M}$}. Then the infinitesimal generator is $\sigma$-strong densely defined. If $\{T_{t}\}$ preserves adjoints it is even $\sigma$-strong* densely defined.
Proof.
This follows from a combination of the previous lemma and Remark 6.3. ∎
Definition 6.7.
[…]
For $T$ as above it is clear that $T$ will be $\sigma$-strong* closable (equivalently $\sigma$-weakly closable by convexity) iff $S(T)=\{0\}$. We use this fact to posit a closability criterion of *-derivations. The proof is based on an idea of Niknam [39]. Closability criteria of this nature seem to have first been recorded by Bratteli and Robinson (see [8, Theorem 4]) using a quite different proof.
Lemma 6.8.
Let $\delta$ be a $\sigma$-strong* densely defined *-derivation on $\mathcal{M}$}.
  1. (1)
    Then $\overline{S(\delta)}^{w*}$ is a two sided *-ideal.
  2. (2)
    If $\mathcal{M}$} is a factor and if there exists a normal state $\omega$ such that $\omega\circ\delta=0$, then $\delta$ is $\sigma$-weakly closable.
Proof.
The first claim is an easy consequence of the product rule for derivations. Specifically given any $b\in S(\delta)$, $b_{0}\in\mathrm{dom}(\delta)$ and $(a_{\gamma})\subset\mathrm{dom}(\delta)$ with $a_{\gamma}\to 0$ and $\delta(a_{\gamma})\to b$, we have that $a_{\gamma}b_{0},b_{0}a_{\gamma}\in\mathrm{dom}(\delta)$. Moreover both $(b_{0}a_{\delta})$ and $(a_{\gamma}b_{0})$ converge $\sigma$-strong* to 0 with $\delta(b_{0}a_{\gamma})=\delta(b_{0})a_{\gamma}+b_{0}\delta(a_{\gamma})\to b_{0}b$ and similarly $\delta(a_{\gamma}b_{0})\to bb_{0}$. So $S(\delta)$ is an $\mathrm{dom}(\delta)$-ideal. Now let $c\in{\mathcal{M}}$ be given. By the density of $\mathrm{dom}(\delta)$ we may select a net $(b_{\gamma})\subset\mathrm{dom}(\delta)$ converging $\sigma$-strong* to $c$. From what we have just shown, $(bb_{\gamma}),(b_{\gamma}b)\subset S(\delta)$. Since by convexity the $\sigma$-weak and $\sigma$-strong* closures of $S(\delta)$ agree, it follows that $bc,cb\in\overline{S(\delta)}^{w*}$.
For the second claim note that since $\overline{S(\delta)}^{w*}$ is a weak* closed ideal, there must exists a central projection $e$ such that $e{\mathcal{M}}=\overline{S(\delta)}^{w*}$. But since $\mathcal{M}$} is a factor we must have $e\in\{0,{\mathds{1}}\}$. Note that by definition $S(\delta)\subseteq\overline{\mathrm{ran}(\delta)}^{w*}$. (Here we used the fact that by convexity the $\sigma$-weak and $\sigma$-strong* closures of $\mathrm{ran}(\delta)$ agree.) So if there is indeed a normal state such that $\omega\circ\delta=0$, then that state annihilates $\overline{S(\delta)}^{w*}$ which in turn means that $e=0$, or equivalently $S(\delta)=\{0\}$. ∎
6.2. Existence of Quantum Killing Lie derivatives
Remark 6.9.
Let $Z$ be a Killing vector field on $(M,g)$. Given $\mathcal{O}\in\mathcal{K}(M,g)$ we know from Corollary 5.6 that there exists $t_{0}>0$ such that for all $t\in[-t_{0},t_{0}]$ the *-homomorphisms $\pi_{t}^{Z}$ are defined on ${\mathcal{M}}(\mathcal{O})$ and implement the action of the local flow $\Psi_{t}^{Z}$ as described in Proposition 4.10. For a fixed $s$ let $\mathscr{T}_{s}$ be the set of all $\mathcal{O}\in\mathcal{K}(M,g)$ for which the *-homomorphisms $\pi_{t}^{Z}$ are defined on ${\mathcal{M}}(\mathcal{O})$ and implement the action of the local flow $\Psi_{t}^{Z}$ for all $t\in[-s,s]$. Now let $D_{S}$ be $\mathrm{span}\{{\mathcal{M}}(\mathcal{O})\colon\mathcal{O}\in\mathscr{T}_{s}\}$. The continuity of the $\pi_{t}^{Z}$’s ensure that the $D_{s}$’s satisfy the criteria of Definition 6.2.
Theorem 6.10.
Let $(M,g)$ be a globally hyperbolic Lorentzian manifold and let ${\mathcal{M}}(M)$ be as before. If for each Killing vector field $Z$ and each $\mathcal{O}\in\mathcal{K}(M,g)$ there exists a normal state $\omega_{\mathcal{O}}$ on ${\mathcal{M}}(\mathcal{O})$ for which we have that $\omega_{\mathcal{O}}\circ\delta_{Z}=0$, the operator $\delta_{Z}$ is then a $\sigma$-strong* densely defined closable unital *-derivation.
Proof.
Remark 6.9 ensures that we have access to the theory of the previous section. Let ${\mathcal{M}}(\mathcal{O})$ be given. The fact that $\delta_{Z}$ is a derivation follows by applying the product rule to the fact that we will for small enough $t$ have that $\pi_{t}^{Z}(ab)=\pi_{t}^{Z}(a)\pi_{t}^{Z}(b)$. Similarly the fact that $\pi_{t}^{Z}(a^{*})=\pi_{t}^{Z}(a)^{*}$ ensures that $\delta_{Z}$ is a *-derivation. In view of the strong continuity noted above, the fact that $\delta_{Z}$ is $\sigma$-strong* densely defined follows from the results of the previous subsection. Then the results in the previous subsection also ensure that $\delta_{Z}$ is $\sigma$-strong* closable on ${\mathcal{M}}(\mathcal{O})$. ∎
Part II Von Neumann algebraic representations of local algebras
7. The selection of appropriate von Neumann algebraic representations
In this Section we will describe in detail von Neumann algebras related to subsets ${\mathcal{O}}\in{\mathcal{K}}(M,g)$ as well as mappings between them with some emphasis on dynamical maps originating from the Klein-Gordon equation. We start with selecting the appropriate algebras that describe the fields. The key observation which makes this selection possible is that the Weyl algebra constructed from the proper symplectic space encompasses, via affiliation, the fundamental (unbounded) observables for Quantum Field Theory in the curved spacetime. We shall here look specifically at regular (and in particular quasi-free) representations of these algebras. We recall that elements of the Weyl algebra, Weyl operators, are of the form
(7.1) $W(f)=e^{i\phi(f)},$
where $\phi(f)$ is the smeared field operator satisfying the commutation relation
$(\phi(f)\phi(g)-\phi(g)\phi(f))\psi=\sigma(f,g)\psi\mbox{ for all }f,g\in{\mathcal{H}},\psi\in\rm{dom}(\phi(f)),$
and $\mathcal{H}$} is a symplectic space. As was already mentioned in the introduction, the description of a system based on von Neumann algebras includes the algebraic structure of observables, while the GNS construction induced by the given state $\omega$ describes the above structure in the representation reflecting the current state of the system. However, when describing a system with curved space-time, dynamic relations given by the (semi-classical) Einstein equation should be taken into account, cf. [50]. It is assumed that the back-reaction of the quantum field on the spacetime is described by the equation
(7.2) $G_{ab}=8\pi\langle T_{ab}\rangle_{\omega},$
where $T_{ab}$ is the stress-energy tensor and $G_{ab}$ the Einstein tensor. It is important to note that $T_{ab}$ can be expressed as a quadratic form of the (unbounded) field operators.
Based on recent results, we can say that there are two possibilities to enlarge the set of observables for the description of the system in question. It is worth noting that if the set of observables is larger, then the set of states (continuous functionals on that set) consists of more regular states. An illustration of this claim was given in [34].
The first option is to limit ourselves to a special class of states, specifically to the subclass of quasi-free states that satisfy Hadamard’s condition. The second is to use Orlicz space techniques as was described in our recent paper [31]. So in this Section we will limit ourselves to quasi-free states with some emphasis on Hadamard states.
Within the framework of algebraic quantization of field theory, we are not interested in using microlocal analysis to describe states that may be analogues of the vacuum in Minkowski space. As far as states are concerned, we are rather interested in those states whose GNS representation well illustrates the presented theory. Among such states the quasi-free states are of course particularly significant in view of the fact that their GNS representation yield Araki-Woods factors. Within this class the Hadamard states are those which are widely used in QFT in particular due to the fact that the expectation value of the quantum stress energy tensor corresponding to a Hadamard state makes sense. As we shall shortly see, there is no need to search far and wide “good” quasi-free states which well illustrate the theory. Such states in fact emerge from the actual geometry of the underlying Lorentzian manifold. We do, however, in section 9 we point out the existence of a structure, thoroughly described in our paper [31], which characterizes states with well-defined expectation values on a large family of unbounded field observables.
We reiterate that the idea of quasi-free state, introduced by Derek Robinson [43], was suggested by QFT where models of free and generalized free fields are characterized by the property that all their Wightman functions $W_{n}$ are uniquely determined by their two point functions $W_{2}$. This concept is also important in Quantum Statistical Physics, see [7] for details.
We show how for any $\mathcal{O}\in\mathcal{K}(M,g)$ we may construct such a quasi-free state conditioned to $\mathcal{O}$. As before we shall follow the notational convention of [3]. Let $\widetilde{G}_{\mathcal{O}}^{\pm}$ be the adv/ret Green’s operators associated with $\mathcal{O}$ and $\widetilde{G}_{\mathcal{O}}=\widetilde{G}_{\mathcal{O}}^{+}-\widetilde{G}_{\mathcal{O}}^{-}$. A crucial ingredient in the construction of this example is the local boundedness of the Green’s operator noted below:
Proposition 7.1.
[19, Proposition 3.1] For any $\mathcal{O}\in\mathcal{K}(M,g)$, the operator $i\widetilde{G}_{\mathcal{O}}$ is a densely defined symmetric operator on $L^{2}(\mathcal{O},dV)$ with domain $\mathscr{D}(\mathcal{O},E)$ which extends to a bounded operator $i\overline{G}_{\mathcal{O}}$ on $L^{2}(\mathcal{O},dV)$.
Example 7.2.
Given $f,g\in\mathscr{D}(\mathcal{O},E)$, we will by the preceding proposition and Lemma 4.3.5 of [3] have that
$|\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}g\rangle\,dV|^{2}$ $=$ $\left|\int_{M}\langle f+\mathrm{ker}(\overline{G}_{\mathcal{O}}),\overline{G}_{\mathcal{O}}(g)\rangle\,dV\right|$
$\leq$ $\|\overline{G}_{\mathcal{O}}\|.\|f+\mathrm{ker}(\overline{G}_{\mathcal{O}})\|.\|g+\mathrm{ker}(\overline{G}_{\mathcal{O}})\|$
$\leq$ $\|\overline{G}_{\mathcal{O}}\|.\|f+\overline{\mathrm{ker}(\widetilde{G}_{\mathcal{O}})}\|.\|g+\overline{\mathrm{ker}(\widetilde{G}_{\mathcal{O}})}\|$
The final inequality follows from the fact that $\overline{\mathrm{ker}(\widetilde{G}_{\mathcal{O}})}\subset\mathrm{ker}(\overline{G}_{\mathcal{O}})$.
We will extract an inner product satisfying equation (5.30) of [30] from the above and use that to construct a quasi-free state. To do this we will make use of the correspondences described in [30, Proposition 5.2.12]. In our setting the space $\mathrm{SOL}(\mathcal{O})$ referred to by [30] will just be $\mathrm{SOL}(\mathcal{O})=\{\widetilde{G}_{\mathcal{O}}(f)\colon f\in\mathscr{D}(\mathcal{O},E)\}$ and the space $\mathcal{E}(\mathcal{O})$ the quotient space $\mathcal{E}=\mathscr{D}(\mathcal{O},E)/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ (to see this consider equation (5.3) of [30] alongside the definition of this space). If one compares equations (5.4) and (5.12) in [30], the the bilinear form $\tau$ they define on $\mathrm{SOL}$ is just $\tau(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))=\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV$. Writing $[f]$ for the equivalence class $f+\overline{\mathrm{ker}(\widetilde{G}_{\mathcal{O}})}$, it is now an exercise to see that $\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV=\int_{M}\langle x_{f},\widetilde{G}_{\mathcal{O}}(y_{g})\rangle\,dV$ for any $x_{f}\in[f]$, $y_{g}\in[g]$. Similarly $\widetilde{G}_{\mathcal{O}}f=\overline{G}_{\mathcal{O}}[f]$. That means that we may write $\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV=\int_{M}\langle[f],\overline{G}_{\mathcal{O}}[g]\rangle\,dV$. For any $f,g\in\mathscr{D}(\mathcal{O},E)$, it follows from the previously displayed inequality that
$|\tau(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))|^{2}\leq\|\overline{G}_{\mathcal{O}}\|^{2}\|[f]\|^{2}\|[g]\|^{2}$
for all $f],[g]\in\mathcal{E}(\mathcal{O})$. We write $H_{[\cdot]}$ for the real Hilbert space generated by $\mathcal{E}(\mathcal{O})$ equipped with the quotient norm norm inherited from $L^{2}(\mathcal{O},dV)/\mathrm{ker}(\overline{G}_{\mathcal{O}})$ Here we worked with real Hilbert spaces for the sake of simplicity, but all of the above can of course be complexified in the obvious manner, which we now assume to be the case. Thus the prescription
$\mu^{\mathcal{O}}([f],[g])=\tfrac{\|\overline{G}_{\mathcal{O}}\|}{2}\langle[f],[g]\rangle\quad f,g\in\mathscr{D}(\mathcal{O},E)$
defines an inner product on $\mathcal{E}$ which will satisfy
$\frac{1}{4}|\tau(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))|^{2}\leq\mu^{\mathcal{O}}([f],[f]).\mu^{\mathcal{O}}([g],[g])\mbox{ for all }[f],[g]\in\mathcal{E}.$
What we need in order to have access to [30, Propositions 5.2.23 & 5.2.24] is much the same thing, but for $\mathrm{SOL}(\mathcal{O})$ not $\mathcal{E}(\mathcal{O})$. Since these two spaces are bijective by [30, Proposition 5.2.12] we may now define a matching inner product $\widetilde{\mu}^{\mathcal{O}}$ on $\mathrm{SOL}(\mathcal{O})$ by simply setting $\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))=\mu^{\mathcal{O}}([f],[g])$ for all $f],[g]\in\mathcal{E}$. That will then furnish us with an inner product which satisfies (5.30) of [30]. By definition (see also Theorem 5.2.24 in [30]) there is then a unique quasi-free state $\omega_{\mathcal{O}}$ corresponding to the two-point functions $(\omega_{\mathcal{O}})_{2}(f,g)=\frac{1}{4}\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)+i\tau(\widetilde{G}_{\mathcal{O}}(f),\widetilde{G}_{\mathcal{O}}(g))$. This follows, see [36], from the recipe for a quasi-free state described there, namely
(7.3) $\omega_{s}(W(x))=\exp\{-\tfrac{1}{2}s(x,x)\}$
where the quadratic form $s(\cdot,\cdot)$ satisfies $|\sigma(f,g)|^{2}\leq s(f,f)s(g,g)$, $\sigma$ is the symplectic form, and there is the corresponding formula for the two-point function. We shall write ${\mathcal{M}}(\mathcal{O})$ for the resultant von Neumann algebra.
7.1. The behaviour of $\{{\mathcal{M}}(\mathcal{O}):\mathcal{O}\in\mathcal{K}(M,g)\}$ with respect to local isometries
Let $\mathcal{O}\in\mathcal{K}(M,g)$ be as before and let $\Psi$ be an isometry which preserves time-orientation. Consider the quasi-free state on ${\mathcal{M}}(\mathcal{O})$ with the two point function given by $(f,g)\to\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)+i\tau(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)$ where $f,g\in\mathscr{D}(\mathcal{O},E)$. We know from Theorem 5.3 that for $f,g\in\mathscr{D}(\mathcal{O},E)$ we have
$\tau(\widetilde{G}_{\mathcal{O}}(g),\widetilde{G}_{\mathcal{O}}(f))$ $=$ $\int_{M}\langle g,\widetilde{G}_{\mathcal{O}}(f)\rangle\,dV$
$=$ $\int_{M}\langle g\circ\Psi^{-1},\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\rangle\,dV$
$=$ $\tau(\widetilde{G}_{\Psi(\mathcal{O})}(g\circ\Psi^{-1}),\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})).$
We show that we similarly have that
$\widetilde{\mu}^{\mathcal{O}}(\widetilde{G}_{\mathcal{O}}f,\widetilde{G}_{\mathcal{O}}g)=\widetilde{\mu}^{\Psi(\mathcal{O})}(\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1}),\widetilde{G}_{\Psi(\mathcal{O})}(g\circ\Psi^{-1})).$
To see this we firstly note that the facts noted about the preservation of measure in the first part of the proof of Theorem 5.3, ensure that
$\|f\|_{\mathcal{O}}^{2}=\int_{M}\langle f,f\rangle\,dV=\int_{M}\langle f\circ\Psi^{-1},f\circ\Psi^{-1}\rangle\,dV=\|f\circ\Psi^{-1}\|^{2}_{\Psi(\mathcal{O})}$
for all $f,g\in\mathscr{D}(\mathcal{O},E)$. When combined with Lemma 5.2, this in fact ensures that that $\|\overline{G}_{\mathcal{O}}\|=\|\overline{G}_{\Psi(\mathcal{O})}\|$. The claimed equality will then follow if we can additionally show that $\langle[f],[g]\rangle=\langle[f\circ\Psi^{-1}],[g\circ\Psi^{-1}]\rangle$ for all $f,g\in\mathscr{D}(\mathcal{O},E)$. This in turn firstly follows from the fact noted above that $f\to f\circ\Psi^{-1}$ defines a unitary from $L^{2}(\mathcal{O},dV)$ to $L^{2}(\Psi(\mathcal{O}),dV)$, and secondly from the fact that by Lemma 5.2 this prescription also bijectively maps $\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ onto $\mathrm{ker}(\widetilde{G}_{\Psi(\mathcal{O})})$. We therefore have that $\|[f]\|_{\mathcal{O}}^{2}=\|[f\circ\Psi^{-1}]\|^{2}_{\Psi(\mathcal{O})}$, as was required.
We proceed to sum up the above analysis. The context is set by the symplectic space, $V(M,\sigma_{{\mathcal{O}}})=\mathscr{D}(M,{\mathcal{O}})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ with the symplectic form $\sigma_{{\mathcal{O}}}(f,g)=\langle\langle g,\widetilde{G}_{\mathcal{O}}(f)\rangle\rangle$ (See [3],(page 129).) Within this context we observed the following.
  1. (1)
    The map $S:f\mapsto f\circ\Psi^{-1}$ : $V(M,\sigma_{\Psi({\mathcal{O}})})\to V(M,\sigma_{{\mathcal{O}}})$ is a symplectic map, i.e. $\sigma_{\Psi({\mathcal{O}})}(Sf,Sg)\equiv\langle\langle g\circ\Psi^{-1}),\widetilde{G}_{\Psi(\mathcal{O})}(f\circ\Psi^{-1})\rangle\rangle=\langle\langle g,\widetilde{G}_{\mathcal{O}}(f)\rangle\rangle=\sigma_{{\mathcal{O}}}(f,g)$, for the $\mathrm{CCR}$ algebras $({\mathcal{A}}({\mathcal{O}}),\sigma_{{\mathcal{O}}})$ and $({\mathcal{A}}(\Psi({\mathcal{O}})),\sigma_{\Psi({\mathcal{O}}}))$ where ${\mathcal{A}}({\mathcal{O}})$ is the Weyl algebra over the symplectic space $V(M,\sigma_{{\mathcal{O}}})$, etc.
  2. (2)
    On ${\mathcal{A}}(({\mathcal{O}}))$ and ${\mathcal{A}}(\Psi({\mathcal{O}}))$ there exist quasi-free states $\omega_{{\mathcal{O}}}$ and $\omega_{\Psi({\mathcal{O}})}$ with the corresponding two-point functions.
  3. (3)
    Denote by $\pi_{\omega_{{\mathcal{O}}}}$ ($\pi_{\omega_{\Psi({\mathcal{O}}}}$) the GNS representation of ${\mathcal{A}}({\mathcal{O}})$ (${\mathcal{A}}(\Psi({\mathcal{O}}))$ induced by the state $\omega_{{\mathcal{O}}}$ ($\omega_{\Psi({\mathcal{O}})}$ respectively). It is worth noting that that as $\mathrm{CCR}$ algebras are simple, both representations have trivial kernel, that is the kernels are equal to $\{0\}$.
  4. (4)
    Consequently, there is a -isomorphism $\alpha^{\pi}:\pi_{\omega_{{\mathcal{O}}}}({\mathcal{A}}({{\mathcal{O}}}))\to\pi_{\omega_{\Psi({\mathcal{O}})}}({\mathcal{A}}(\Psi({\mathcal{O}}))$ such that $\omega_{\Psi({\mathcal{O}})}\circ\alpha^{\pi}=\omega_{{\mathcal{O}}}$. It is an easy observation that the last equality leads to: $\alpha^{\pi}(\cdot)=V^{*}\cdot V$ where $V$ is the surjective isometry such that $V:{\mathcal{H}}_{\omega_{\Psi({\mathcal{O}})}}\to{\mathcal{H}}_{\omega_{{\mathcal{O}}}}$, and ${\mathcal{H}}_{\omega_{\Psi({\mathcal{O}})}}$ and ${\mathcal{H}}_{\omega_{{\mathcal{O}}}}$ the corresponding Hilbert spaces from the GNS construction.
We therefore arrive at the following:
Proposition 7.3.
$\alpha^{\pi}$ has a weak extension $\alpha^{\pi}_{ext}$ which is the -isomorphism between ${\mathcal{M}}({\mathcal{O}})$ and ${\mathcal{M}}(\Psi({\mathcal{O}}))$. In other words, there are well-defined mappings between the nets of von Neumann algebras ${\mathcal{O}}\mapsto{\mathcal{M}}({\mathcal{O}})$ and $\Psi({\mathcal{O}})\mapsto{\mathcal{M}}(\Psi({\mathcal{O}}))$ where ${\mathcal{O}}\in{\mathcal{K}}(M,g)$.
7.2. Isotonic properties of $\{{\mathcal{M}}(\mathcal{O}):\mathcal{O}\in\mathcal{K}(M,g)\}$
Our task here is essentially to compare $\widetilde{\mu}_{1}$ and $\widetilde{\mu}_{2}$. At the risk of a clash of notation we will denote the $L^{2}$-closure of a subspace $E$ of $L^{2}(\mathcal{O}_{k},dV)$ by $[E]$. Here $H_{\mu_{k}}$ is of course then a scaled version of $H_{[\cdot]_{k}}=[\mathscr{D}(\mathcal{O},E)]/[\mathrm{ker}(\widetilde{G}_{k})]$. To complete this task we require the following proposition. In proving this proposition we shall constantly make use of the fact that the PDO $P_{M}$ corresponding to the Klein-Gordon equation will map $C^{\infty}(\mathcal{O},E)$ back $C^{\infty}(\mathcal{O},E)$ and that the operator $P_{\mathcal{O}}$ is nothing but the restriction of $P_{M}$. (See [3, Proposition 3.5.1] and the discussion surrounding it.)
Proposition 7.4.
Let $\mathcal{O}\in\mathcal{K}(M,g)$ and $f\in C^{\infty}(\mathcal{O},E)$ be given. If $f$ has bounded partial derivatives of all orders, then there exists $(f_{n})\subset\mathscr{D}(\mathcal{O},E)$ such that $P_{\mathcal{O}}f_{n}\to P_{\mathcal{O}}f$ in $L^{2}(\mathcal{O},dV)$.
Proof.
Step 1: We will first prove the proposition for the case where the manifold in question is $\mathbb{R}^{N}$. We prove that there exists a sequence of functions $(f_{n})\subset\mathscr{D}(\mathcal{O},E)$ which converges to $f$ inside the Sobolev space $W^{2,2}(\mathcal{O})$. Since on $\mathbb{R}^{N}$ the Klein-Gordon operator is a PDO with a very regular structure, it will then follow from this fact and the theory of Sobolev spaces that we must then have that $P_{\mathcal{O}}f_{n}\to P_{\mathcal{O}}f$ in $L^{2}(\mathcal{O},d\lambda_{N})$.
We firstly note that the assumption on $f$ ensures that we will for any multi-index $\alpha$ with $|\alpha|\leq 2$ have that $\int_{\mathcal{O}}\langle D^{\alpha}(f),D^{\alpha}(f)\rangle\,d\lambda_{N}\leq\|D^{\alpha}(f)\|_{\infty}^{2}\lambda_{N}(\mathcal{O})<\infty$ (where $\lambda_{N}$ is Lebesgue measure). So $f$ clearly belongs to the Sobolev space $W^{2,2}(\mathcal{O})$. We now introduce the open subsets $\mathcal{O}_{k}=\{x\in\mathcal{O}:d(x,\partial\mathcal{O})\leq 2^{-k}\}$. We clearly have that $\mathcal{O}_{k}\subset\mathcal{O}_{k+1}$ for all $k$ with $\mathcal{O}=\cup_{k=1}\mathcal{O}_{k}$. A similar argument to what we used above now shows that we will here for any multi-index $\alpha$ with $|\alpha|\leq 2$ have that $\int_{\mathcal{O}\backslash\mathcal{O}_{k}}\langle D^{\alpha}(f),D^{\alpha}(f)\rangle\,d\lambda_{N}\leq\|D^{\alpha}(f)\|_{\infty}^{2}\lambda_{N}(\mathcal{O}\backslash\mathcal{O}_{k})\to 0$. That ensures that for the Sobolev space $W^{2,2}(\mathcal{O})$ we have that $\|f\chi_{\mathcal{O}\backslash\mathcal{O}_{k}}\|_{2,2}\to 0$ as $k\to\infty$.
We may also for any $k$ select a bump function $\zeta_{k}$ such that $\zeta_{k}=1$ on $\overline{\mathcal{O}_{k}}$ and $\zeta_{k}=0$ on $\mathcal{O}\backslash\overline{\mathcal{O}_{k+1}}$ (see the corollary to [51, Theorem 1.11]). The function $\zeta_{k}f$ will then be a function of compact support which agrees with $f$ on $\mathcal{O}_{k}$ and which is supported on $\mathcal{O}_{k+1}$. We therefore have that
$f-\zeta_{k}.f=f\chi_{\mathcal{O}\backslash\mathcal{O}_{k}}-\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}.$
We already know that $\|f\chi_{\mathcal{O}\backslash\mathcal{O}_{k}}\|_{2,2}\to 0$ as $k\to\infty$. So if we can show that inside $W^{2,2}(\mathcal{O})$ the function $\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}$ is the limit of a sequence of smooth functions of compact support, it will therefore follow that inside $W^{2,2}(\mathcal{O})$, the function $f$ is then similarly the limit of a sequence of smooth functions of compact support. To prove that this the case, we use the theory of mollifiers (see subsection C.4 in [18]). Given such a mollifier $\eta_{\epsilon}$ it is an exercise to see that we will for any locally integrable function $g$ on $\mathcal{O}$ have that $(\eta_{\epsilon}*g)(x)=\int_{B(0,\epsilon)}\eta_{\epsilon}(y)g(x-y)\,dy=0$ whenever $d(x,\mathrm{supp}(g))>\epsilon$. For any $\epsilon\geq 2^{-k-2}$ the function $\eta_{\epsilon}*(\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}})$ will therefore be supported on $\overline{\mathcal{O}_{k+2}}\backslash\mathcal{V}_{k}$ where $\mathcal{V}_{k}=\{x\in\mathcal{O}_{k}:d(x,\partial\mathcal{O}_{k})>2^{-k-2}\}$.
Assuming this to be the case, the functions $\eta_{\epsilon}*(\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}})$ are then all smooth functions of compact support [18, §C.4, Theorem 6] for which the restrictions to any open subset $V$ with $\overline{V}\subset\mathcal{O}$ will converge to $\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}$ inside $W^{2,2}(V)$ as $\epsilon\to 0$ [18, §5.3.1, Theorem 1]. But $\mathcal{O}_{k+2}$ is such a subset and each $\eta_{\epsilon}*(\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}})$ as well as $\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}$ are supported on $\mathcal{O}_{k+2}$. Hence as $\epsilon\searrow 0$, the functions $\eta_{\epsilon}*(\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}})$ even converge to $\zeta_{k}.f\chi_{\mathcal{O}_{k+1}\backslash\mathcal{O}_{k}}$ in $W^{2,2}(\mathcal{O})$. This then proves the proposition in the $\mathbb{R}^{N}$ case.
Step 2: The manifold $(M,g)$ admits an orientable atlas, and so by compactness $\overline{\mathcal{O}}$ can be covered with finitely many oriented charts $U_{1},\dots,U_{m}$ (see [40, Definition 1.41]) which correspond to relatively compact open regions in $\mathbb{R}^{N}$ via diffeomorphisms $\rho_{i}$ from $V_{i}$ to $U_{i}$. We can also arrange matters so that the Jacobian $|\mathrm{det}(d\rho_{i})|$ is bounded. (To see this note that for such triple $(U,V,\rho)$ where $U$ is a neighbourhood of $p\in\mathcal{O}$ we can first replace $U$ with a smaller compact neighbourhood $K$ (ensuring the compactness of $\rho_{i}^{-1}(K)$ and the boundedness of the Jacobian, and then replace $U$ with the interior of $K$.) The compact set $\overline{\mathcal{O}}$ can then be covered with such modified charts, with compactness ensuring that only finitely many such charts are necessary. For any given $U_{i}$ we will only work with $U_{i}\cap\mathcal{O}$. In so doing we will wherever convenient assume that functions supported on $U_{i}\cap\mathcal{O}$ are extended to all of $U_{i}$ by assigning the value 0 on $U_{i}\backslash(U_{i}\cap\mathcal{O})$ (and similarly for functions supported on $\rho_{i}^{-1}(U_{i}\cap\mathcal{O}$). Since $f\in C^{\infty}(\mathcal{O},E)$ is a function with partial derivatives of all orders bounded, the function $f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i}$ is similarly a smooth function with partial derivatives of all orders bounded. Writing $\widetilde{P}_{i}$ for the coordinatisation of $P_{U_{i}}$, it follows from step 1 that there exists a sequence $(g_{n,i})\subset\mathscr{D}(U_{i}\cap\mathcal{O},E)$ for which $(\widetilde{P}_{i}(g_{n,i}))$ converges to $\widetilde{P}_{i}(f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i})$ in $L^{2}(\rho_{i}^{-1}(U_{i}\cap\mathcal{O}),\lambda_{N})$-norm. The diffeomorphisms $\rho_{i}$ now induce measurable isomorphisms between the Borel algebras $\mathcal{B}(U_{i})$ and $\mathcal{B}(V_{i})$ with the Jacobian playing the role of a Radon-Nikodym derivative (see [24, Prop 5.19]). We therefore have that
$\int_{U_{i}\cap\mathcal{O}}\langle P_{i}(f|_{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1}),P_{i}(f|_{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1})\rangle\,dV$
$=$ $\int|\langle\widetilde{P}_{i}(f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n,i}),\widetilde{P}_{i}(f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n,i})\rangle|\,|\mathrm{det}(d\rho_{i})|\,d\lambda_{N}$
$\leq$ $\|\,|\mathrm{det}(d\rho_{i})|\,\|_{\infty}\int|\langle\widetilde{P}_{i}(f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n_{i}}),\widetilde{P}_{i}(f|_{U_{i}\cap\mathcal{O}}\circ\rho_{i})-\widetilde{P}_{i}(g_{n,i})\rangle\,d\lambda_{N}$
[…]
$\int_{M}\langle P_{\mathcal{O}}(f-f_{n}),P_{\mathcal{O}}(f-f_{n})\rangle\,dV$
$\leq$ $\sum_{i=1}^{m}\int_{U_{i}\cap\mathcal{O}}\langle P_{i}(f|_{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1}),P_{i}(f|_{U_{i}\cap\mathcal{O}})-P_{i}(g_{n,i}\circ\rho_{i}^{-1})\rangle\,dV$
$\to$ $0\mbox{ as }n\to\infty$
Corollary 7.5.
For any $f\in\mathrm{ker}(\widetilde{G}_{2})$, the restriction $f|_{\mathcal{O}_{1}}$ will belong to $\mathrm{ker}(\widetilde{G}_{1})]$.
Proof.
We remind the reader that $\mathrm{ker}(\widetilde{G}_{k})=P_{\mathcal{O}_{k}}(\mathscr{D}(\mathcal{O}_{k},E))$. (See page 129 of [3].) So let $f=P_{\mathcal{O}_{2}}(g)$, where $g\in\mathscr{D}(\mathcal{O}_{2},E)$, be given. (Here $P_{\mathcal{O}}$ is the partial differential operator corresponding to the Klein-Gordon equation on $\mathcal{O}$.) The corollary will then follows from the preceding proposition once we notice that since it is the restriction of a function of compact support, the function $g|_{\mathcal{O}_{1}}\in C^{\infty}(\mathcal{O}_{1},E)$ must have bounded partial derivatives of all orders. ∎
Proposition 7.6.
Let $e_{(2:1)}$ be the projection from $H_{[\cdot]_{2}}$ onto the closure of $\{[f_{ext}]_{2}:f\in\mathscr{D}(\mathcal{O}_{1},E)\}$. Then the map
$U:f_{ext}+\overline{\mathrm{ker}(\widetilde{G}_{1})}\mapsto f+\overline{\mathrm{ker}(\widetilde{G}_{1})}\quad f\in\mathscr{D}(\mathcal{O}_{1},E)$
extends to a unitary operator from $e_{(2:1)}H_{[\cdot]_{2}}$ onto $H_{[\cdot]_{1}}$.
Proof.
Given $f\in\mathscr{D}(\mathcal{O}_{1},E)$ with $\widetilde{G}_{1}(f)\neq 0$, let $b_{f}\in[\mathrm{ker}(\widetilde{G}_{1})]$ be the unique element for which $f-b_{f}\perp[\mathrm{ker}(\widetilde{G}_{1})]$. Select $(b_{n})\subset\mathscr{D}(\mathcal{O}_{1},E)$ such that $P_{\mathcal{O}_{1}}(b_{n})\to b_{f}$. (Here we are using the fact noted earlier that $\mathrm{ker}(\widetilde{G}_{k})=P_{\mathcal{O}_{k}}(\mathscr{D}(\mathcal{O}_{k},E))$.) Since (again as noted earlier) $P_{\mathcal{O}_{2}}(b_{n}^{ext})=P_{\mathcal{O}_{1}}(b_{n})_{ext}$ for all $n$, we have that $(P_{\mathcal{O}_{1}}(b_{n})_{ext})\subset P_{\mathcal{O}_{2}}(\mathscr{D}(M_{\mathcal{O}_{2}},E))=\mathrm{ker}(\widetilde{G}_{2})$. The fact that $\lim_{n}P_{\mathcal{O}_{1}}(b_{n})_{ext}=b_{f}^{ext}$, therefore ensures that $b_{f}^{ext}\in[\mathrm{ker}(\widetilde{G}_{2})]$. Given any $\upsilon\in\mathrm{ker}(\widetilde{G}_{2})$, we may now use the preceding corollary and the fact that $f$ and $b_{f}$ are supported on $\mathcal{O}_{1}$, to see that
$\int_{M}\langle(f-b_{f})_{ext},\upsilon\rangle,dV=\int_{M}\langle f-b_{f},\upsilon|_{\mathcal{O}_{1}}\rangle,dV=0.$
Thus $(b_{f})_{ext}$ is the unique element of $\mathrm{ker}(\widetilde{G}_{2})]$ for which $(f-b_{f})_{ext}\perp[\mathrm{ker}(\widetilde{G}_{2})]$. But that means that $\|f+[\mathrm{ker}(\widetilde{G}_{1})]\|=\|f-b_{f}\|=\|f_{ext}+[P_{\mathcal{O}_{2}}(C_{0}^{\infty}(M_{\mathcal{O}_{2}},E)]]_{2}\|$. Thus the canonical map from $ext(\mathscr{D}(\mathcal{O}_{1},E))+\mathrm{ker}(\widetilde{G}_{2}]/[\mathrm{ker}(\widetilde{G}_{2})]$ to $\mathscr{D}(\mathcal{O}_{1},E))]/[\mathrm{ker}(\widetilde{G}_{1})]$ is in fact a surjective linear isometry. ∎
By suitably rescaling the above unitary operator, we arrive at the following
Corollary 7.7.
There is a unitary operator $v_{(2:1)}$ from $e_{(2:1)}H_{\mu_{2}}$ to $H_{\mu_{1}}$ which maps each $f_{ext}+\overline{\mathrm{ker}(\widetilde{G}_{2})}$, where $f\in\mathscr{D}(\mathcal{O}_{1},E)$, to $f+\overline{\mathrm{ker}(\widetilde{G}_{1})}$.
We now pass to comparing the algebras ${\mathcal{M}}(\mathcal{O}_{2},\mathcal{O}_{1})$ and ${\mathcal{M}}(\mathcal{O}_{1})$. By [30, Theorem 5.2.24] the GNS-spaces $H_{\omega_{k}}$ ($k=1,2$) are constructed from the action of the field operators on the Fock spaces $\mathcal{F}^{+}(H_{\widetilde{\mu}_{k}})$. We leave it as an exercise to verify that $e_{(2:1)}$ lifts to a projection $\Gamma(e_{(2:1)})$ from $\mathcal{F}^{+}(H_{\mu_{2}})$ onto $\mathcal{F}^{+}(e_{(2:1)}H_{\mu_{2}})$, and $v_{(2:1)}$ to a unitary operator $\Gamma(v_{(2:1)})$ from $\mathcal{F}^{+}(e_{(2:1)}H_{\mu_{2}})$ to $\mathcal{F}^{+}(H_{\mu_{1}})$, where $\Gamma(\cdot)$ stands for the second quantization map.
Let ${\mathcal{M}}(\mathcal{O}_{2}:\mathcal{O}_{1})$ denote the subalgebra of ${\mathcal{M}}(\mathcal{O}_{2})$ generated by the Weyl operators with symbols in $\mathrm{SOL}(\mathcal{O}_{2}:\mathcal{O}_{1})$ where $\mathrm{SOL}(\mathcal{O}_{2}:\mathcal{O}_{1})$ stands for the subset $\{\widetilde{G}_{2}(f_{ext}):f\in\mathscr{D}(\mathcal{O}_{1},E)\}$ of $\mathrm{SOL}(\mathcal{O}_{2})$. We claim that ${\mathcal{M}}(\mathcal{O}_{2}:\mathcal{O}_{1})$ commutes with $\Gamma(e_{(2:1)})$. The representation of $\mathrm{CCR}(\mathcal{O}_{2})$ on $\mathcal{F}^{+}(H_{\mu_{2}})$, therefore compresses to a representation of of $\mathrm{CCR}(\mathcal{O}_{2}:\mathcal{O}_{1})$ on $\mathcal{F}^{+}(e_{(2:1)}H_{\mu_{2}})$. To see this notice that for any $f\in\mathscr{D}(\mathcal{O}_{1},E)$ and $x_{1},\dots,x_{n}\in\mathrm{SOL}(\mathcal{O}_{2})$ we have that
$e_{(2:1)}^{n+1}(\widetilde{G}_{2}(f_{ext}))\otimes x_{1}\otimes\dots\otimes x_{n}$ $=$ $\widetilde{G}_{2}(f_{ext}))\otimes e_{(2:1)}(x_{1})\otimes\dots\otimes e_{(2:1)}(x_{n})$
$=$ $\widetilde{G}_{2}(f_{ext}))\otimes e_{(2:1)}^{n}(x_{1}\otimes\dots\otimes x_{n})$
and hence that $\Gamma(e_{(2:1)})a^{*}(\widetilde{G}_{2}(f_{ext}))=a^{*}(\widetilde{G}_{2}(f))\Gamma(e_{(2:1)})$. By duality we then similarly have that $\Gamma(e_{(2:1)})a(\widetilde{G}_{2}(f_{ext}))=a(\widetilde{G}_{2}(f_{ext}))\Gamma(e_{(2:1)})$, and hence that $\Gamma(e_{(2:1)})\phi(\widetilde{G}_{2}(f_{ext}))=\phi(\widetilde{G}_{2}(f_{ext}))\Gamma(e_{(2:1)})$. The Borel functional calculus now ensures that $\Gamma(e_{(2:1)})\pi_{\omega_{2}}(W(\widetilde{G}_{2}(f_{ext})))=\pi_{\omega_{2}}(W(\widetilde{G}_{2}(f_{ext})))\Gamma(e_{(2:1)}).$
Careful checking will similarly show that for any $f\in\mathscr{D}(\mathcal{O}_{1},E)$ we will in terms of $\Gamma(v_{(2:1)})$ have that $\Gamma(v_{(2:1)})a^{*}(\widetilde{G}_{2}(f_{ext}))=a^{*}(\widetilde{G}_{1}(f))\Gamma(v_{(2:1)})$ and hence that $\Gamma(v_{(2:1)})a^{*}(\widetilde{G}_{2}(f_{ext}))\Gamma(v_{(2:1)})^{*}=a^{*}(\widetilde{G}_{1}(f))$. Taking adjoints shows that we also have that $\Gamma(v_{(2:1)})a(\widetilde{G}_{2}(f_{ext}))\Gamma(v_{(2:1)})^{*}=a(\widetilde{G}_{1}(f))$, and so the same will be true of the field operators, namely that $\Gamma(v_{(2:1)})\phi(\widetilde{G}_{2}(f_{ext}))\Gamma(v_{(2:1)})^{*}=\phi(\widetilde{G}_{1}(f))$. So by the Borel functional calculus we must surely have that
$\Gamma(v_{(2:1)})\pi_{\omega_{2}}(W(\widetilde{G}_{2}(f_{ext})))\Gamma(v_{(2:1)})^{*}=\pi_{\omega_{1}}(W(\widetilde{G}_{1}(f))).$
The pair $({\mathcal{M}}(\mathcal{O}_{2},\mathcal{O}_{1}),\mathcal{F}^{+}(e_{(2:1)}H_{\mu_{2}}))$ is therefore unitarily isomorphic to $({\mathcal{M}}(\mathcal{O}_{1}),\mathcal{F}^{+}(H_{\mu_{1}}))$. That then yields the following observation:
Proposition 7.8.
The pair $({\mathcal{M}}(\mathcal{O}_{2},\mathcal{O}_{1}),\mathcal{F}^{+}(e_{(2:1)}H_{\mu_{2}}))$ is unitarily isomorphic to the pair $({\mathcal{M}}(\mathcal{O}_{1}),\mathcal{F}^{+}(H_{\mu_{1}}))$.
7.3. Conclusion
A comparison of the previous two subsections now yield the following conclusion:
Theorem 7.9.
With respect to the *-isomorphisms and compressions described above, the collection of pairs $\{({\mathcal{M}}(\mathcal{O}),H_{\omega_{\mathcal{O}}})):\mathcal{O}\in\mathcal{K}(M,g)\}$ is an example of the category VN-STDFM introduced in section 3, and hence describes a locally covariant theory in the sense of section 3 of [20].
Since the Hadamard states are a subset of quasi-free states, the above conclusion is obviously valid for these states. It is well known (see [50]) that Hadamard states have well-defined expectation values of the stress-energy tensor. Since the description of Hadamard states would go beyond the scope of this work we will limit ourselves to only a few remarks. The key property of these states is given in the language of distribution theory. And this description is well adapted to the Wightman axioms (also called Gårding-Wightman axioms), i.e. field operators localized at a point of space-time. In this paper, we analyze quantum fields localized in space-time regions. Therefore, continuing the strategy described in the paper [31], we will in the next section provide conditions also leading to the description of (regular) quantum fields with a well defined average value of the stress-energy tensor operator.
8. States invariant under the action of a Killing vector field $Z$
Let $(M,g)$ be a globally hyperbolic spacetime which is stationary, i.e. $(M,g)$ admits a time-like Killing field $Z$ as well as a one parameter group of isometries $\beta_{t}^{Z}:M\to M$ whose orbits are time-like. Consider the Weyl algebra ${\mathcal{A}}(M)$ associated to a real scalar Klein-Gordon field as was described in [30]. It was shown in [30] that there exists the quasi-free state $\omega_{Z}$ on ${\mathcal{A}}(M)$ which is invariant under the action of $Z$, i.e. $\omega_{Z}\circ\pi^{Z}_{t}=\omega_{Z}$ where $\pi^{Z}_{t}(W(\psi))=W(\psi\circ\beta^{Z}_{t}))$; cf. Section 5.2.7 in [30].
Consequently, taking the GNS representation $\pi_{\omega_{z}}(\cdot)$ of ${\mathcal{A}}(M)$ induced by the state $\omega_{Z}$, see [30], we arrive at the dynamical system
(8.1) $({\mathfrak{M}}(M),\alpha^{Z}_{t}),$
where ${\mathfrak{M}}(M)$ is the weak closure of $\pi_{\omega_{Z}}({\mathcal{A}}(M))$, $\alpha^{Z}_{t}(\pi_{\omega_{Z}}(a))=U^{(Z)}_{t}\pi_{\omega_{Z}}(a){U_{t}^{{(Z)}}}^{*}=\pi_{\omega_{Z}}(\pi^{Z}_{t}(a))$, for all $t\in{\rm I\!R}$ and $a\in{\mathcal{A}}(M)$.
Strong continuity $t\mapsto U^{(Z)}_{t}$ of the unitary group $\{U^{(Z)}_{t}\}$ ensures the existence of an infinitesimal generator which may then be interpreted as the natural Hamiltonian for the considered system. We pause to justify the claim of continuity. As is pointed out in the discussion following equation (5.40) in [30], strong continuity of this unitary group is equivalent to the requirement that
(8.2) $\lim_{t\to 0}\omega_{Z}(a^{*}\alpha^{Z}_{t}(a))=\omega_{Z}(a^{*}a)\mbox{ for all }a\in{\mathcal{A}}(M).$
This will of course follow if the automorphism group $\pi_{\omega_{Z}}(a)\to\pi_{\omega_{Z}}(\alpha^{Z}_{t}(a))$ where $a\in{\mathcal{A}}(M)$ is strongly continuous at 0. With $\mu_{Z}$ denoting the inner product that forms the real part of the two point function of $\omega_{Z}$ (see [30, Theorem 5.2.24]), we claim that this automorphism group will indeed exhibit such strong continuity if for all $f\in\mathscr{D}(M,E)$ we have that $\mu_{Z}(G(f)-G(f)\circ\beta_{t}^{Z},G(f)-G(f)\circ\beta_{t}^{Z})\to 0$ as $t\to 0$. Right at the outset we note that the isometries $\beta_{t}^{Z}$ commute with the Green’s operator $G$. (See for example page 226 of [15].) This ensures that the action of these isometries on $\mathscr{D}(M,E)$, lift to an action on $\{G(f)\colon f\in\mathscr{D}(M,E)\}$. We shall constantly make silent use of this fact. Since ${\mathcal{A}}(M)$ is generated by the Weyl operators and since the span of vectors of the form $\pi_{\omega_{Z}}(W(G(f)))\Omega_{\omega_{Z}}$ is dense in the GNS Hilbert space, this claim will follow if we can show that for any $f,g\in\mathscr{D}(M,E)$ we have that $\|\pi_{\omega}((W(G(f)\circ\beta^{Z}_{t})-W(G(f)))W(G(g)))\Omega_{\omega}\|_{\omega}\to 0$ as $t\to 0$. Here we dropped the subscript $Z$ for the sake of simplicity. In the following we also set $\psi=G(f)$ and $\varphi=G(g)$ for the sake of simplicity. Now observe that
$\|\pi_{\omega}((W(\psi\circ\beta^{Z}_{t})-W(\psi)))W(\varphi))\Omega_{\omega}\|^{2}_{\omega}$
$\quad=\|\pi_{\omega}((W(\psi)^{*}W(\psi\circ\beta^{Z}_{t})-{\mathds{1}}))W(\varphi))\Omega_{\omega}\|^{2}_{\omega}$
$\quad=\|\pi_{\omega}((W(-\psi)W(\psi\circ\beta^{Z}_{t})-{\mathds{1}}))W(\varphi))\Omega_{\omega}\|^{2}_{\omega}$
$\quad=\|\pi_{\omega}((e^{i\sigma(\psi,\psi\circ\beta_{t}^{Z})/2}W(\psi\circ\beta^{Z}_{t}-\psi)-{\mathds{1}}))W(\varphi))\Omega_{\omega}\|^{2}_{\omega}$
$\quad=\omega(2{\mathds{1}}-e^{i\sigma(\psi,\psi\circ\beta_{t}^{Z})/2}W(\varphi)^{*}W(\psi\circ\beta^{Z}_{t}-\psi)W(\varphi)$-
$\quad\qquad e^{-i\sigma(\psi,\psi\circ\beta_{t}^{Z})/2}W(\varphi)^{*}W(\psi\circ\beta^{Z}_{t}-\psi)^{*}W(\varphi))$
$\quad=2-e^{i\sigma(\psi,\psi\circ\beta_{t}^{Z})/2}e^{i\sigma(\varphi,\psi\circ\beta_{t}^{Z}-\psi)}\omega_{Z}(W(\psi\circ\beta^{Z}_{t}-\psi))$-
$\quad\qquad e^{i\sigma(\psi,\psi\circ\beta_{t}^{Z})/2}e^{i\sigma(\varphi,\psi-\psi\circ\beta_{t}^{Z})}\omega_{Z}(W(\psi-\psi\circ\beta^{Z}_{t}))$
$\quad=2-e^{i\sigma(\psi+2\varphi,\psi\circ\beta_{t}^{Z}-\psi)/2}\omega_{Z}(W(\psi\circ\beta^{Z}_{t}-\psi))$-
$\quad\qquad e^{i\sigma(2\varphi-\psi,\psi-\psi\circ\beta_{t}^{Z})/2}\omega_{Z}(W(\psi-\psi\circ\beta^{Z}_{t}))$
$\quad=2-e^{-i\sigma(2\varphi+\psi,\psi\circ\beta_{t}^{Z}-\psi)/2}e^{-\mu_{Z}(\psi-\psi\circ\beta^{Z}_{t},\psi-\psi\circ\beta^{Z}_{t})/2}$-
$\quad\qquad e^{i\sigma(2\varphi-\psi,\psi-\psi\circ\beta_{t}^{Z})/2}e^{-\mu_{Z}(\psi-\psi\circ\beta^{Z}_{t},\psi-\psi\circ\beta^{Z}_{t})/2}$
(In the third last equality we used the fact that $W(\varphi)^{*}W(\psi)W(\varphi)=e^{-i\sigma(\varphi,\psi)}W(\psi)$ - a fact which can easily be verified from the properties of Weyl operators noted in Remark 5.1.) The symplectic form $\sigma$ we have to work with here is of course $\sigma(G(f),G(g))=\int_{M}\langle f,G(g)\rangle\,dV$. For this form we do have that
$\sigma(G(f)\circ\beta^{Z}_{t}(f),G(2g\pm f))=\int_{M}\langle f\circ\beta_{t}^{Z},G(2g\pm f)\rangle\,dV\to$
$\int_{M}\langle f,G(2g\pm f)\rangle\,dV=\sigma(G(f),G(2g\pm f))\mbox{ as }t\to 0.$
So $\sigma(G(2g\pm f),G(f)\circ\beta_{t}^{Z}-G(f))\to 0$ as $t\to 0$. If we now combine this fact with the previously centred equations, it clearly follows that if for all $f\in\mathscr{D}(M,E)$ we have that $\mu_{Z}(G(f)-G(f)\circ\beta_{t}^{Z},G(f)-G(f)\circ\beta_{t}^{Z})\to 0$ as $t\to 0$, then $\|\pi_{\omega}((W(f\circ\beta^{Z}_{t})-W(f))W(g))\Omega_{\omega}\|_{\omega}\to 0$ as $t\to 0$. As noted earlier this is sufficient to ensure that equation (8.2) will hold whenever $\lim_{t\to 0}\mu_{Z}(G(f)-G(f)\circ\beta_{t}^{Z},G(f)-G(f)\circ\beta_{t}^{Z})=0$ for all $f\in\mathscr{D}(M,E)$.
Remark 8.1.
It is an easy observation that the dynamical system $({\mathfrak{M}}(M),\alpha^{Z}_{t})$ given above provides a nice illustration of the Section 5. Namely, taking derivatives of $\alpha^{(Z)}_{t}$ will lead to Quantum Killing Lie derivatives. Here it is the invariance $\omega_{Z}\circ\pi^{Z}_{t}=\omega_{Z}$ noted above that ensures that Theorem 6.10 is applicable thereby ensuring the existence of these derivatives as densely defined closable *-derivations.
9. Enlarging the observable algebras
Turning to the question of enlarging the set of observables so that the enlarged set allows for the existence of arbitrary moments of the field operators, we note that $\pi_{\omega_{Z}}(W(f))=e^{i\phi_{\omega_{Z}}(f)}$, where $\phi_{\omega_{Z}}(f)$ are smeared self-adjoint field operators. It is worth noting that, by the construction, $\{\phi_{\omega_{Z}}(f)\}$ are affiliated to ${\mathfrak{M}}(M)$.
We recall, that the Orlicz space technique applied to regular fields satisfying an $H$-boundedness condition with respect to the density of the dual weight of the reference weight $\omega$ of the von Neumann algebra, leads to the space $L^{\cosh-1}({\mathfrak{M}})$. For all details see the paper [31]. The key observation to be made here, is that under such an $H$-boundedness restriction, the field operators affiliated to ${\mathfrak{M}}(M)$ which have all moments well defined may in a very natural way be embedded into this space. In other words, if the field operators $\{\phi_{\omega_{Z}(f)}\}$ satisfy an $H$-boundedness condition, then they can be canonically embedded into $L^{\cosh-1}({\mathfrak{M}}(M))$. In that case $\langle T_{ab}\rangle_{\omega_{Z}}$ exists and consequently the (semiclassical) Einstein equation would be well defined. In situations where we have a strongly continuous action $\alpha_{t}$ that not only leaves $\omega$ invariant but also corresponds to the modular automorphism group of some other fns weight, one can use results like [49, Corollary VIII.3.6] and [4, Proposition 4.2] to show that in the crossed product ${\mathfrak{M}}\rtimes_{\omega}\mathbb{R}$ this action is implemented by a unitary group with generator a product of the density of the dual weight and an operator affiliated to the centralizer of $\mathfrak{M}$}.
It is crucial to note that the Orlicz space technique allows for the consideration of a broader class of states than quasi-free states. Namely, the natural states for $L^{cosh-1}({\mathfrak{M}}(M))$-observables are normalized positive functionals corresponding to elements in the space $L\log(L+1)({\mathfrak{M}})$, see [35], [31]. The above indicates that the proposed description of regular fields operators can be used to analyze even (strongly) interacting fields having strong quantum correlations, for example entanglements, see Subsection 3.2.
We end this section with the following remarks on $H$-boundedness. In the present paper we are concerned with showing the usefulness of von Neumann algebras in the description of quantum systems in curved space-time. In particular, the AQFT framework was formulated in this language. When discussing Hadamard states in the previous section we mentioned Wightman’s axioms. It is key to note that quantum fields described by Wightman axiomatics and satisfying the $H$-boundedness condition can be equally described in the AQFT language. More precisely, there is a correspondence between quantum fields in the sense of Wightman with the test function space belonging to the theory of ultradistribution and regular field operators affiliated to the net of local algebras, see [21], [29], [46]. In this work we were concerned with the net of local von Neumann algebras and field operators affiliated to such nets. Therefore, the $H$-bounded condition can be considered a natural requirement.
Part III Conclusion: the emergent framework and future development
10. Concluding remarks
In this work, which is a continuation and an extension of our recent paper [31], we, firstly, present the AQFT framework in terms of von Neumann algebras. The presented framework is natural from a categorical point of view. Furthermore, the mentioned approach allows us to formulate the framework for AQFT without at the outset assuming the existence of a global Hilbert space. The important point to note here is that the presented approach allows the description of the entanglement phenomenon in QFT. In the next part of the paper, which is devoted to the study of differential calculus within the framework of field dynamics, we introduced the description of flows on local algebras without using the concept of “tangentially conditioned” algebras. In this way, we obtain a complementary and more simplified description of flows than that in our previous paper [31]. Here, flows for local quantum systems in curved space-time are defined directly using Killing vector fields acting on spacetime. The basic properties of quantum local flows have been shown. Furthermore, the conditions under which quantum Lie derivatives exist are provided. This was achieved by studying quantum local flows on regions of spacetime. We also showed the one may use the geometry of the underlying Lorentzian manifold itself to construct very natural quasi-free states at a local level, with the resultant net of local algebras satisfying the criteria of the category VN-STDFM, which in turn ensures that this net is an appropriate framework for a locally covariant quantum field theory. Hence the natural remaining question is whether the additivity property for quantum fields holds in the proposed description of nets of local algebras. For a deeper discussion of additivity the reader is refered to [44], [47]. We will here only note that the additivity property was very useful for the description of a quantum flows given in Theorem 4.1 in [31]. It is worth pointing out that for a quasi-free state satisfying mild regularity conditions, the additivity property will in the present setting also hold. To see this we need the concept of a partition of unity.
Definition 10.1.
A collection of subsets $\mathcal{U}=\{u_{\alpha}\colon\alpha\in A\}$ of a manifold $M$ is called locally finite, if for all $p\in M$ there is an neighborhood $V_{p}$ of $p$ with $V_{p}\cap U_{\alpha}$ non-empty for only finitely many of the sets $U_{\alpha}$.
Definition 10.2.
A partition of unity on a manifold $M$ is a collection $\{x_{\gamma}\colon\gamma\in\Gamma\}$ of smooth real-valued functions such that
  1. (1)
    $\{\mathrm{supp}(x_{\gamma})\colon\gamma\in\Gamma\}$ is locally finite,
  2. (2)
    $x_{\gamma}(p)\geq 0$ for all $p\in M$ and all $\gamma\in\Gamma$,
  3. (3)
    $\sum_{\gamma\in\Gamma}x_{\gamma}(p)=1$ for all $p\in M$
If $\mathcal{U}=\{U_{\alpha}\colon\alpha\in A\}$ is an open cover of $M$ we say that a partition of unity is subordinate to the open cover $\mathcal{U}$ if for every $\gamma$ there is an $\alpha$ such that $\mathrm{supp}(x_{\gamma})\subseteq U_{\alpha}$.
The key fact we need is the following existence theorem.
Theorem 10.3.
[51, Theorem 1.11] Let $M$ be a $d$-dimensional $C^{k}$-differentiable second countable manifold and $\mathcal{W}$ any open cover. Then $M$ admits a countable partition of unity subordinate to the cover $\mathcal{W}$ with the support of each function in the t of unity compact.
Theorem 10.4.
Let $\mathcal{O}\in\mathcal{K}(M,g)$ be given and let $\omega$ be a quasi-free state on ${\mathcal{A}}(\mathcal{O})$ of the form described in [30, Theorem 5.2.24] for which the inner product $\mu$ respects the mode of convergence introduced in [3, Definition 4.3.6]. Then there exists a countable family $\mathcal{V}_{n}$ of globally hyperbolic charts in $\mathcal{O}$ such that $\cup_{n}\mathcal{V}_{n}=\mathcal{O}$. Moreover ${\mathcal{M}}(\mathcal{O})=(\cup_{n=1}^{\infty}{\mathcal{M}}(\mathcal{V}_{n}))^{\prime\prime}$ where ${\mathcal{M}}(\mathcal{O})$ is the double commutant of $\pi_{\omega}({\mathcal{A}}(\mathcal{O}))$ and each ${\mathcal{M}}(\mathcal{V}_{n})$ is the double commutant of the copy of ${\mathcal{A}}(\mathcal{V}_{n})$ inside $\pi_{\omega}({\mathcal{A}}(\mathcal{O}))$.
Proof.
In the latter part of the proof we will silently make extensive use of the ideas and notation used in Section 8. A degree of familiarity with the flow of ideas in Section 8 is therefore necessary to navigate the proof. For each $p\in\mathcal{O}$ there exists a chart $\mathcal{V}_{p}\subset\mathcal{O}$ of $p$ which belongs to $\mathcal{K}(M,g)$ (see [37, Theorem 2.7]). We therefore have that $\cup_{p}\mathcal{V}_{p}=\mathcal{O}$. We know from Moretti’s notes [38, Definition 8.2] that $\mathcal{O}$ is second countable and hence that there must exist a countable subfamily $\{\mathcal{V}_{n}\colon n\in\mathbb{N}\}$ which still covers $\mathcal{O}$. For these charts it is clear that $(\cup_{n=1}^{\infty}{\mathcal{M}}(\mathcal{V}_{n}))^{\prime\prime}\subseteq{\mathcal{M}}(\mathcal{O})$. We are therefore left with the challenge of proving the converse containment. It is here that we need the concept of a partition of unity. Aside from what we noted earlier about second countability, it is also clear from Moretti’s notes (see the discussion following [38, Definition 2.23]) that globally hyperbolic Lorentzian manifolds satisfy all the prerequisites of Theorem 10.3. Hence there exists a countable partition of unity $\{x_{n}\colon n\in\mathbb{N}\}$ on $\mathcal{O}$ with compact supports subordinate to the family $\{\mathcal{V}_{n}\colon n\in\mathbb{N}\}$.
For the next part of the proof we need to understand in which manner the $\mathrm{CCR}$ $C^{*}$-algebras ${\mathcal{A}}(\mathcal{V}_{n})$ live “inside” ${\mathcal{A}}(\mathcal{O})$. Given $f,g\in\mathscr{D}(\mathcal{V}_{n},E)$ and denoting the extensions to $\mathcal{O}$ by $f_{ext}$, $g_{ext}$, the fact that the supports are in $\mathcal{V}_{n}$ ensures that
$\int_{M}\langle f,\widetilde{G}_{\mathcal{V}_{n}}(g)\rangle\,dV=\int_{M}\langle f_{ext},\widetilde{G}_{\mathcal{O}}(g_{ext})\rangle\,dV.$
(Here we used the fact that $\widetilde{G}_{\mathcal{V}_{n}}(g)=\widetilde{G}_{\mathcal{O}}(g_{ext})_{|\mathcal{V}_{n}}$.) The left and right hand side of the displayed equation above each respectively represent degenerate symplectic forms on the spaces $\mathcal{E}(\mathcal{V}_{n})=\{f\colon f\in\mathscr{D}(\mathcal{V}_{n},E)\}$ and $\mathcal{E}(\mathcal{O}:\mathcal{V}_{n})\{a\colon a\in\mathscr{D}(\mathcal{O},E),\mathrm{supp}(a)\subset\mathcal{V}_{n}\}$. By [3, Lemma 4.3.8] the map $\widetilde{G}_{\mathcal{V}_{n}}(f)\to\widetilde{G}_{\mathcal{O}}(f_{ext})$ will map $\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ into $\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ and hence induces a well-defined map from $\mathcal{E}(\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ to the subspace $\mathcal{E}(\mathcal{O}:\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ of $\mathcal{E}(\mathcal{O})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$. We remind the reader that the spaces $\mathcal{E}(\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ and $\mathcal{E}(\mathcal{O})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ are respectively canonically bijective to $\mathrm{SOL}(\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{V}_{n}}(f)\colon f\in\mathscr{D}(\mathcal{V}_{n},E)\}$ and $\mathrm{SOL}(\mathcal{O}:\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{O}}(a)\colon a\in\mathscr{D}(\mathcal{O},E),\mathrm{supp}(f)\subset\mathcal{V}_{n}\}$. In view of the equality in the previously displayed equation it is now clear that induced map from $\mathcal{E}(\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{V}_{n}})$ to the subspace $\mathcal{E}(\mathcal{O}:\mathcal{V}_{n})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ of $\mathcal{E}(\mathcal{O})/\mathrm{ker}(\widetilde{G}_{\mathcal{O}})$ is a symplectic isomorphism between these two spaces. By means of the aforementioned bijective correspondences, it may of course equivalently be regarded as a symplectic isomorphism from $\mathrm{SOL}(\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{V}_{n}}(f)\colon f\in\mathscr{D}(\mathcal{V}_{n},E)\}$ to $\mathrm{SOL}(\mathcal{O}:\mathcal{V}_{n})=\{\widetilde{G}_{\mathcal{O}}(f)\colon f\in\mathscr{D}(\mathcal{O},E)\}$. Hence the $\mathrm{CCR}$ $C^{*}$-algebras generated by these two sets are *-isomorphic. So when regarded as a subalgebra of ${\mathcal{A}}(\mathcal{O})$, we may identify ${\mathcal{A}}(\mathcal{V}_{n})$ with the $\mathrm{CCR}$ algebra generated by $\mathrm{SOL}(\mathcal{O}:\mathcal{V}_{n})$. In the appropriate representation we then take the double commutants of these algebras to get the corresponding statement for von Neumann algebras.
Given any $f\in\mathscr{D}(\mathcal{O},E)$, we now set $f_{n}=f.x_{n}$. Checking shows that $f=\sum_{n=1}^{\infty}f_{n}$. Given any $f_{n}$, we know from the preceding theorem that there exists some $k\in\mathbb{N}$ with $\mathrm{supp}(f_{n})\subset\mathcal{V}_{k}$. Checking now shows that $f_{n}=(g_{n})_{ext}$ where $g_{n}=(f_{n})_{|\mathcal{V}_{k}}$. It is clear that $g_{n}$ belongs to $\mathscr{D}(\mathcal{V}_{k},E)$, and hence that the Weyl operator $W(\widetilde{G}_{\mathcal{O}}(f_{n}))$ belongs to ${\mathcal{M}}(\mathcal{V}_{k})$. It is also clear from the properties of the Weyl operators that the Weyl operator $W(\sum_{n=1}^{k}f_{n})$ is a just a multiple of the product $\Pi_{n=1}^{k}W(f_{n})$. To conclude we therefore need to show that in the representation engendered by $\omega$, $W(f)$ is the $\sigma$-strong limit of the sequence $(W(\sum_{n=1}^{k}f_{n}))_{k=1}^{\infty}$.
For any compact subset $K$ of $M$, the local finiteness of the partition of unity ensures that for any $p\in K\cap\mathrm{supp}(f)$ we can find a neighbourhood $U_{p}\subset\mathcal{O}$ of $p$ with $U_{p}\cap\mathrm{supp}(f_{n})$ non-empty for at most finitely many of $f_{n}$’s. By compactness we may cover $K\cap\mathrm{supp}(f)$ with finitely many such neighbourhoods, from which it then follows that there must exist some $N_{K}\in\mathbb{N}$ such that $(\sum_{n=1}^{k}f_{n})_{|K}=f_{|K}$ for all $k\geq N_{K}$. A similar argument shows that the same is true for any of the derivatives of the $f$ and the partial sums $\sum_{n=1}^{k}f_{n}$. This of course means $(\sum_{n=1}^{k}f_{n})_{k=1}^{\infty}$ and each matching sequence of derivatives converges uniformly to $f$ each of its corresponding derivatives on any compact subset of $\mathcal{O}$. This has two consequences:
Firstly for any $g\in\mathrm{SOL}(\mathcal{O}:\mathcal{V}_{n})$ we have that $\int_{M}\langle\sum_{n=1}^{k}f_{n},\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV\to\int_{M}\langle f,\widetilde{G}_{\mathcal{O}}(g)\rangle\,dV$ as $k\to\infty$. To see this note that integration with respect to the volume form can be shown to make sense for $C_{0}(M)$ (the continuous functions vanishing at infinity). On appealing to the Riesz representation theorem it now follows that integration with respect to the volume form may be regarded as integration with respect to a Radon measure. Since $\mathcal{O}$ has finite measure with respect to this Radon measure (due to the compactness of its closure), the claim now follows from basic facts regarding Radon measures and what we noted above. Secondly since $f$ and each $\sum_{n=1}^{k}f_{n}$ is supported on the compact set $\mathrm{supp}(f)$, it follows that here $(\sum_{n=1}^{k}f_{n})_{k=1}^{\infty}$ converges to $f$ in the mode described in [3, Definition 3.4.6]. But by [3, Proposition 3.4.8], $(\widetilde{G}_{\mathcal{O}}(\sum_{n=1}^{k}f_{n}))_{k=1}^{\infty}$ will then converge to $\widetilde{G}_{\mathcal{O}}(f)$ in exactly the same mode. If therefore on $\mathrm{SOL}(\mathcal{O})$ the inner product $\mu$ respects this mode of convergence, we will have that $\mu(\widetilde{G}_{\mathcal{O}}(f)-\widetilde{G}_{\mathcal{O}}(\sum_{n=1}^{k}f_{n}),\widetilde{G}_{\mathcal{O}}(f)-\widetilde{G}_{\mathcal{O}}(\sum_{n=1}^{k}f_{n}))\to 0$ as $k\to\infty$. The on replacing $G(f)\circ\beta_{t}^{Z}$ with $\widetilde{G}_{\mathcal{O}}(\sum_{n=1}^{k}f_{n})$, the same type of argument as was used in Section 8 now shows that in the representation engendered by $\omega$, $W(f)$ is, as required, the $\sigma$-strong limit of the sequence $(W(\sum_{n=1}^{k}f_{n}))_{k=1}^{\infty}$. ∎
All that makes it legitimate to form
Conjecture 10.5.
Let $\mathcal{O}\in\mathcal{K}(M,g)$. Write $\mathrm{QuantLie}$ for the set
$\{\delta_{Z}\colon Z\mbox{ a Killing vector field}\}.$
Then ${\mathcal{M}}^{\infty}(\mathcal{O},\mathrm{QuantLie})$ is weak* dense in ${\mathcal{M}}(\mathcal{O})$.
If the conjecture is true, one can use the ideas of du Bois-Violette [16] to develop non-commutative differential geometric structures for QFT, c.f. the discussion in [31].